### v3

# GORAG: Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

Abstract.
Text classification is a fundamental task in data mining, pivotal to various applications such as tabular understanding and recommendation. Although neural network-based models, such as CNN and BERT, have demonstrated remarkable performance in text classification, their effectiveness heavily relies on abundant labeled training data. This dependency makes these models less effective in dynamic few-shot text classification, where labeled data is scarce, and new target labels frequently appear based on application needs. Recently, large language models (LLMs) have shown promise due to their extensive pretraining and contextual understanding ability. Current approaches provide LLMs with text inputs, candidate labels, and additional side information (e.g., descriptions) to classify texts. However, their effectiveness is hindered by the increased input size and the noise introduced through side information processing. To address these limitations, we propose a graph-based online retrieval-augmented generation framework, namely GORAG, for dynamic few-shot text classification. Rather than treating each input independently, GORAG constructs and maintains a weighted graph by extracting side information across all target texts. In this graph, text keywords and labels are represented as nodes, with edges indicating the correlations between them. To model these correlations, GORAG employs an edge weighting mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a minimum-cost spanning tree tailored for each text input. Empirical evaluations demonstrate that GORAG outperforms existing approaches by providing more comprehensive and precise contextual information.
Large language models, online retrieval-augmented generation, few-shot learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/02
Figure 1. An example of the DFSTC task with two rounds. In round 1, an example text is classified as Machine Learning. In round 2, an example text is classified as Computer Vision.
1. Introduction
Text classification is a fundamental task of text data mining and is connected to various tasks, such as tabular understanding (Chen et al., 2024c; Wang et al., 2024c; Wen et al., 2024a, b; He et al., 2025; Zheng et al., 2023), and recommendation (Du et al., 2024a; Han et al., 2024b; Wang et al., 2024b; Li et al., 2023; Zheng et al., 2024; Chiang et al., 2023; Du et al., 2024b). In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Prieto et al., 2016; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters et al., 2018; Wang et al., 2018; Yang et al., 2019; Miao et al., 2021), such as CNN (Kim, 2014), Bert (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on text classification tasks. However, the effectiveness of these NN-based approaches depends on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018). Furthermore, in real-world applications, such as Web Of Science (Kowsari et al., 2017) and IMDb (imd, [n.d.]), target labels for text often change based on the application’s requirements (Xia et al., 2021), leading to the dynamic few-shot text classification (DFSTC) task. As shown in Figure 1, the DFSTC task begins with a few initial classes (e.g., Machine Learning). As new topics like Computer Vision emerge in later rounds, models must adapt to these changes and accurately classify new examples with minimal labeled data. Hence, how to develop methods to dynamically classify text with limited labeled data available remains a necessity and open problem.
Depending the technique for the DFSTC task, current models can be mainly categorized into two types, i.e., data augmentation-based models and RAG-based models. Firstly, data augmentation-based models (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021) create additional training data by mixing the pairs of the few-shot labeled data and assigning mixed labels indicating the validity for these created data based on the labels of each data pairs (Xia et al., 2021). Also, several data augmentation approaches create extra semantic-related content based on the label names (Meng et al., 2018; Meng et al., 2020) to enrich the training corpus further. However, due to the limited labeled data, and the label names may not always be available, the generated text data can have very limited patterns. Consequently, the text classification models trained on these generated data may over-fit on limited text data and are not generalizable (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023).
Figure 2. An overview of GraphRAG (Edge et al., 2024). It first extracts and indexes entities and relations as graph nodes and edges. Then it performs Top-K similarity-based retrieval for query-related subgraphs. Finally, it classifies texts based on the subgraph summary and the query text.
Recently, large language models (LLMs) (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), pre-trained on extensive corpora, have achieved significant success due to their superior and comprehensive text understanding abilities. However, researchers (Gao et al., 2023; Peng et al., 2024) discover that LLMs can struggle to understand their inputs due to a lack of task-specific knowledge. Therefore, the Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024) were proposed to combine the predicted text, candidate labels, and the retrieved side information (e.g., descriptions of text and labels, or documents) as LLM inputs, enable LLMs to understand the predicted text and labels more comprehensively with the help of side information from external sources. However, the incorporation of side information can further increase the input size and noise (Chen et al., 2023), which impedes the efficiency and effectiveness of LLMs (Sarthi et al., 2024).
As a result, Compression-based RAG approaches (Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024) are proposed to compress the length of context and extract the key information. Specifically, LongLLMLingua (Xia et al., 2021) and CompaAct (Yoon et al., 2024) propose filtering uninformative texts based on the query. However, these approaches may result in information loss and degrade overall performance. Then, to address this, graph RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024; Peng et al., 2024; Zhang et al., 2025; Han et al., 2024a), such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2024), propose to to index the unstructured texts into the structured graphs. For example, in Figure 2, the basic idea of graph RAG considers text entities as graph nodes, and text sentences describing relations between these entity nodes as graph edges, later, it retrieves subgraphs base on the query text, and generates short summaries of these subgraphs as the side information for LLMs, reducing the input length for LLMs. However, existing Compression RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) still have three issues in the DFSTC task.
  1. (1)
    Uniform-indexing issue: Some of these approaches construct graphs by indexing extracted text chunks uniformly. They do not consider the varying importance and extraction confidence of each text chunk, which may provide incorrect and unreliable context for LLMs.
  2. (2)
    Threshold-dependent issue: These approaches select relevant information for each input text based on a globally predefined threshold. However, the optimal retrieval threshold can vary across different text samples; hence, relying on a globally predefined threshold can be suboptimal for the entire dataset.
  3. (3)
    Narrow-source issue: These approaches only retrieve side information from the each few-shot training text and ignore the important information among different the query texts, consequently limiting their retrieve accuracy and comprehensiveness.
To address these issues, we propose a novel hyperparameter-free Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot text classification, called GORAG. In general, GORAG constructs and maintains an adaptive online weighted graph by extracting side information from all target texts, and tailoring the graph retrieval for each input.
Firstly, GORAG constructs a weighted graph with keywords extracted from labeled texts as keyword nodes, and text labels as label nodes. The edge of this graph represents the relationship between keywords and labels. Secondly, GORAG employs an edge-weighting mechanism to assign different weights to edges created by the last step. An example of the constructed weighted graph is shown in Figure 3. The edge weights represent the keywords’ importance and relevance to the respective text’s label. Thirdly, GORAG constructs a minimum-cost spanning tree on the constructed weighted graph based on keywords of the query text, and then retrieves the label nodes within the spanning tree as candidate labels. Since the generated spanning tree is determined solely by the constructed graph and the keywords of the query text, GORAG achieves adaptive retrieval without relying on any human-defined retrieval thresholds, making the retrieval more precise regarding each query text. Lastly, GORAG applies an online indexing mechanism, which can enrich the weighted graph with the keywords from query texts, enhancing its few-shot learning performance by providing a more comprehensive retrieve source.
We summarize our novel contributions as follows.
  • We present a hyperparameter-free online RAG framework for DFSTC tasks, namely GORAG. The GORAG framework consists of three steps: graph construction, candidate label retrieval, and text classification, aim at addressing the uniform-index, threshold-dependent and narrow-source issues.
  • We develop a novel graph edge weighting mechanism based on the text keywords’ importance within the text corpus, which enables our approach to effectively model the relevance between keywords and labels, thereby helping the more accurate retrieval.
  • To avoid any human-selected thresholds, we formulate the candidate label retrieval problem which is akin to the NP-hard Steiner Tree problem. To solve this problem, provide an efficient and effective solution to generate candidate labels.
2. Preliminary and Related Works
We first introduce the preliminaries of dynamic few-shot text classification in Section 2.1 and then discuss the related works for DFSTC task in Section 2.2. The important notations used in this paper are listed in Table 6 in Appendix A.
2.1. Dynamic Few-shot Text Classification Task
Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world application that involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$. Traditional approaches require fine-tuning with large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available. Recently, few-shot learning addresses this limitation by enabling models to classify text with only a small number of labeled examples per class (Meng et al., 2018; Meng et al., 2020). On the other hand, dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024d) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones. Combining these aspects, Dynamic Few-Shot Text Classification (DFSTC) task allows the model to handle evolving classification tasks with minimal labeled data, and better suited for real world scnerio (Xia et al., 2021).
In DFSTC task, the model is provided with multiple rounds of new class updates. Specifically, In each round $r$, a new set of classes $\mathcal{Y}_{new}^{r}$ is introduced, and the labelled dataset for $\mathcal{Y}_{new}^{r}$ is denoted as $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{new}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y_{i}\in\mathcal{Y}_{new}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$. Also, we denote candidate cumulative labels from the first round to the $r$-th round as $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}_{new}^{i}$. Formally, at the $r$-th round, given the candidate labels $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}_{new}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}={f}^{r}_{\theta}(t_{u},\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{u}$. Then, we can get the predicted label $y^{r*}\in\mathcal{Y}^{r}$ for the unseen text $t_{u}$ as follows.
(1) $y^{r*}=\arg\max_{y^{r}\in\mathcal{Y}^{r}}({f}_{\theta}^{r}(t_{u},\mathcal{Y}^{r})[y^{r}])$
DFSTC is valuable for its ability to learn from limited labeled data, adapt to evolving class distributions, and address real-world scenarios where classes and data evolve over time.
2.2. Dynamic Few-shot Text Classification Model
Current DFSTC models can be broadly categorized into NN-based models and RAG-based models, RAG-based models can be further categorized into Long Context RAG models, and Compression-based RAG models.
2.2.1. NN-based models
NN-based models (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021) generate additional data contrastively based on the few-shot labeled data and the text formed label names to train the NN-based classifier. However, due to the limited labeled data, the generated text data of these models can have very limited patterns, which makes them prone to overfitting (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023). Also, as the text-formed label names are not always available in real-world scenarios, these NN-based models may not always be applicable.
Figure 3. An example of the constructed weighted graph, where red nodes denote the label nodes and black nodes denote the keyword nodes. The lower the edge cost is, the more related the keyword is to the respective label.
Figure 4. An overview of GORAG in each round. In Part 1, GORAG constructs a weighted graph based on keywords extracted from the few-shot training data. In Part 2, GORAG performs adaptive graph retrieval and outputs the candidate labels, which is a subset of the original target label set. In Part 3, GORAG first classifies texts into one of the candidate labels, and then applies online indexing to update the graph with newly extracted nodes from query text.
2.2.2. Long Context RAG models
Recently, Large Language Model (LLM)-based models have undergone rapid development (Fernandez et al., 2023; Amer-Yahia et al., 2023; Zhang et al., 2024c; Miao et al., 2024; Snyder et al., 2024; Zhong et al., 2024; Yan et al., 2024; Chen et al., 2024b; Ding et al., 2024) and have been successfully adapted to various data mining tasks, including semantic parsing (Drozdov et al., 2022; Mekala et al., [n.d.]; Zhong et al., 2024; Lei et al., 2024; Sinha et al., 2024; Zhou et al., 2024), spatial-temporal mining (Zhang et al., 2024d; Li et al., 2024b; Yuan et al., 2024; Liu et al., 2024b; Wang et al., 2024a; Ning and Liu, 2024), and graph mining (Chen et al., 2024b; Tang et al., 2024; Cai et al., 2024; Jiang et al., 2024b; Cheng et al., 2024; Chen et al., 2024a; Fang et al., 2024). Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), making them originally suitable for dynamic text classification tasks. However, the lack of fine-tuning can lead LLMs to generate incorrect answers, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023). To mitigate hallucinations, researchers try to retrieve text documents outside of LLM as side information to help the LLM inference, namely Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Hu et al., 2024; Wu et al., 2024a; Cai et al., 2024; Oosterhuis et al., 2024; Yadav et al., 2024; Che et al., 2024). However, the retrieved contents from these models remain unstructured and can be lengthy, which impedes the efficiency and effectiveness of LLMs, leading to lost-in-the-middle issues (Liu et al., 2024a).
2.2.3. Compression-based RAG models
To address the issue suffered by Long Context RAG models, compression-based RAG models were proposed (Peng et al., 2024; Zhang et al., 2024b; Wu et al., 2024b; Li et al., 2024a; Zhang et al., 2024a; Jiang et al., 2024a; Guo et al., 2024; Gutiérrez et al., 2024), these models try to compress the side information to reduce the LLM input length. Based on how they compress model inputs, these models can be classified into prompt compressor models, and graph-based RAG models. On the one hand, prompt compressor models (Pan et al., 2024; Jiang et al., 2024a), apply LLM’s generation perplexity on to filter out un-important tokens in the model input. However, these models retrieve side information from each document individually, ignoring the inter-document correlations (Edge et al., 2024; Guo et al., 2024). On the other hand, graph-based RAG models (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) convert text chunks from the side information document as graph nodes, and the relation between these text chunks as graph edges. Based on the pre-defined retrieval threshold, they then retrieve a limited number of graph nodes and edges to represent the long document. However, the graph-based RAG models usually consider each graph edge uniformly and ignore the varying confidence and importance of relations between text chunks. Also, as the optimal retrieval threshold can vary across different data samples, their globally fixed threshold selected by humans may be suboptimal for the entire dataset.
In this paper, we proposed a novel compression-based graph-RAG model, namely GORAG, which first constructs a weighted graph to model the inter-document correlations and varying confidence and importance of relations, then it achieves adaptive retrieval on this graph, avoiding any human-defined thresholds.
3. Methodology
3.1. Framework Overview
To address the aforementioned uniform-indexing, threshold-dependent, and narrow-source issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph. As shown in Figure 4, GORAG consists of three core parts, i.e., index graph construction, graph retrieval, and text classification and online Index.
Part 1: Graph Construction. It targets to construct the weighted graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ based on the labeled data $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ at the $r$-th round. An example the weighted graph created on WOS dataset is shown in Figure 3, graph formed like this will be used to provide retrieved-augmented information as context for query texts, enabling LLMs to better understand query texts and accurately predict their labels. Specifically, under each new data $\mathcal{D}_{new}^{r}$ at the $r$-th round, we first extract keywords $\mathcal{K}_{new}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$ from the labeled training texts $\mathcal{T}^{r}_{new}=\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$. Then, we will assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}_{new}^{r}$ and their respective text’s label $y\in\mathcal{Y}^{r}_{new}$. We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance with in the text corpus and relatedness to the label $y$. At last, we merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$. More details can refer to Section 3.2.
Part 2: Graph Retrieval. The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the query text $t$ to the constructed weighted graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all these keywords. From this minimum-cost spanning tree (MST), the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$. As this MST solely depends on the graph itself and the extracted keywords, we can get rid of any human-defined thresholds and achieve adaptive retrieval. For more details, please refer to Section 3.3.
Part 3: Text Classification and Online Index. In Text Classification part, we first classify the query text with a LLM. This process create LLM input with the query text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the weighted graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$ if text formed label names are available. The LLM will carry out classification with this input. After classification, the online indexing procedure dynamically indexes keywords, denoted as $\mathcal{V}_{\text{notexist}}^{t}$, which are extracted from the query text $t$ but are not in the existing graph $\mathcal{G}^{r}$. We then integrate these keywords into the graph to further enrich its structure, with edge weights assigned based on their importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the query text $t$. It enhances the model’s ability to make more accurate predictions. For more details, please refer to Section 3.4.
3.2. Part 1: Graph Construction
In this subsection, we introduce the graph construction procedure of GORAG, which includes a novel edge-weighting mechanism to address the uniform indexing issue. The pseudo code of graph construction is shown in Algorithm 1 in Appendix Appendix C.
For each round $r$, given its respective labeled training texts $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}_{new}$ from $\mathcal{D}_{new}^{r}$. Specifically, it uses the LLM model with an extraction instruction prompt $p_{extract}=$Please extract some keywords from the following passage” for this extaction. Then, we can get all text keywords $\mathcal{K}^{r}_{new}$ based on texts $\mathcal{D}^{r}_{new}$ as follows.
(2) $\mathcal{K}^{r}_{new}=\bigcup_{t\in\mathcal{D}^{r}_{new}}LLM(p_{extract},t).\vspace{-5px}$
These keywords are served as graph nodes, we denote them as keyword nodes.
Also, GORAG incorporats new candidate labels $\mathcal{Y}^{r}_{new}$ at the $r$-th round as graph nodes, denoted as label nodes. Hence, the graph node set $\mathcal{V}^{r}_{new}$ at the $r$-th round can be obtained as follows.
(3) $\mathcal{V}_{new}^{r}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}.$
Then, GORAG add edges between each keyword node $v\in\mathcal{K}_{{new}}^{r}$ to its corresponding label node $y\in\mathcal{Y}_{{new}}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
Next, considering the keywords $\mathcal{K}_{new}^{r}$ from the labeled text are not uniformly related to the text’s label $\mathcal{Y}_{new}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label edge. This weight can be regarded as the correlation between keywords and each label. Firstly, we apply the normalized TF-IDF score (Sammut and Webb, 2010) as our correlation score $CS(v,t)$ to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}_{new}^{r}$ w.r.t. the respective text $t$ it extracted from:
(4) $\vspace{-5px}CS(v,t)=\frac{count(v,t)}{|t|}\times log\frac{|\mathcal{T}^{r}|}{1+|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|},$
where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}_{new}^{i}$ denotes all training and query texts seen so far, $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$.
As the keyword $v$ can be extracted from multiple texts with different labels in different rounds, the final edge weight $w_{v,y}^{r}$ of edge $e_{v,y}$ at the $r$-th round is calculated as an average of all weights from all seen texts with label $y$:
(5) $\vspace{-3px}w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,y,v}}1-CS(v,t_{j})}{|\mathcal{T}^{r,y,v}|},$
where $\mathcal{T}^{r,y,v}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text that contains keyword $v$ and labeled $y$. We denote the generated weighted graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where $\mathcal{V}^{r}_{new}$, $\mathcal{E}^{r}_{new}$, and $\mathcal{W}^{r}_{new}$ denotes the node, edge and edge weight set respectively, an example of the generated graph is shown in Figure 3.
Lastly, $\mathcal{G}^{r}_{new}$ will be merged into the graph from previous rounds $\mathcal{G}^{r-1}$ to form the full graph $\mathcal{G}^{r}$ at round $r$:
(6) $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r}),$
$\mathcal{V}^{r}=\mathcal{V}^{r}_{new}\cup\mathcal{V}^{r-1},$ $\ \mathcal{E}^{r}=\mathcal{E}^{r}_{new}\cup\mathcal{E}^{r-1},\ \mathcal{W}^{r}=\mathcal{W}^{r}_{new}\cup\mathcal{W}^{r-1}.$
Particularly, we define $\mathcal{G}^{0}=\emptyset$ and $\ r\geq 1$. Also, to guarantee the graph connectivity of the resulting graph, we add edges between every new label node $y_{n}\in\mathcal{Y}_{new}^{r}$ and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{y_{n},y_{o}}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is defined as the weighted average of all edge weights that link keyword nodes with label node $y_{n}$ or $y_{o}$:
(7) $\vspace{-5px}\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}\vspace{-5px}$
(8) $\vspace{-3px}w_{y_{n},y_{o}}^{r}=\frac{1}{2}\big{(}\frac{\sum_{v\in\mathcal{M}_{y_{n}}^{r}}w_{v,y_{n}}^{r}}{2\times|\mathcal{M}_{y_{n}}^{r}|}+\frac{\sum_{v\in\mathcal{M}_{y_{o}}^{r}}w_{v,y_{o}}^{r}}{2\times|\mathcal{M}_{y_{o}}^{r}|}\big{)}$
where $\mathcal{N}^{r}(y)$ denote the neighbor nodes of label node $y$ in $\mathcal{G}^{r}$.
After merge, the graph $\mathcal{G}^{r}$ would be used for future retrieval, and be further updated by GORAG’s online indexing mechanism.
3.3. Part 2: Graph Retrieval
In this subsection, we introduce the graph retrieval procedure of GORAG. With the graph constructed in Part 1, GORAG can adaptively retrieve a set of candidate class labels with keywords extracted from query texts without any human-defined thresholds, addressing the threshold dependent issue.
To begin with, GORAG extract keywords $\mathcal{V}_{test}^{t}$ for each query text $t\in\mathcal{T}_{test}$ in the same manner with Equation (2), then, $\mathcal{V}_{test}^{t}$ would be splitted into two subsets:
(9) $\mathcal{V}_{exist}^{t}\cup\mathcal{V}_{notexist}^{t}=\mathcal{V}_{test}^{t},\ \mathcal{V}_{exist}^{t}\cap\mathcal{V}_{notexist}^{t}=\emptyset,$
where $\mathcal{V}_{exist}^{t}$ and $\mathcal{V}_{notexist}^{t}$ denotes the keywords in $\mathcal{V}_{test}^{t}$ that already exist and not yet exist in $\mathcal{G}^{r}$ at current round $r$ respectively. Later, $\mathcal{V}_{exist}^{t}$ would be applied for achieving adaptive retrieval, and $\mathcal{V}_{notexist}^{t}$ would be applied for online indexing to further enrich the constructed graph (Further illustrated in Section 3.4).
To achieve the adaptive retrieval, GORAG trys to find the minimum cost spanning tree (MST) that contain all keyword nodes within $\mathcal{V}_{exist}^{t}$, and then retrieve all label nodes within the generated MST as candidate labels. The intuition behind this approach is that a MST spans the entire graph to cover all given nodes with the smallest possible spanning cost. Consequently, label nodes within the generated MST can be considered important for demonstrating the features of the given keyword node set. As the generation of an MST is a classical combinatorial optimization problem with an optimal solution determined solely by the set of given nodes. By generating the MST and retrieving all label nodes within it, we eliminate the need for any human-defined thresholds, achieving adaptive retrieval.
Table 1. Statistics of the WOS, Reuters, and IFS-Rel datasets, where we divide their original labels into multiple rounds. We achieve a balanced testing set on the Reuters and IFS-Rel datasets by assigning each label 10 and 40 testing samples, respectively.
Dataset Avg. Text Token R1 R2 R3 R4 Total
Testing data Label # Testing data Label # Testing data Label # Testing data Label # Testing data
WOS (Kowsari et al., 2017) 200 2,417 32 2,842 53 2,251 30 1,886 18 9,396
Reuters (Bao et al., 2020) 168 80 8 80 8 80 8 70 7 310
IFS-Rel (Xia et al., 2021) 105 640 16 640 16 640 16 640 16 2,560
Definition 3.1 (Adaptive Candidate label Generation Problem).
Given an undirected weighted connected graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$, a set of keywords $\mathcal{V}_{exist}^{t}$ extracted from text $t$ that can be mapped to nodes in $\mathcal{V}^{r}$, and the target label set $\mathcal{Y}^{r}\in\mathcal{V}^{r}$ at the $r$-th round, our target is to find a set of labels nodes ${\mathcal{Y}}^{r}_{t}\in{\mathcal{Y}}^{r}$. Firstly, we identify a subgraph $\mathcal{G}^{r}_{t}(\mathcal{V}^{r}_{t},\mathcal{E}^{r}_{t},\mathcal{W}^{r}_{t})$ of $\mathcal{G}^{r}$ by minimizing the edge weight sum as follows.
$\min_{e^{r}_{u.v}\in\mathcal{E}^{r}_{t}}w^{r}_{u,v}$
(10) $s.t.$ $v\in\mathcal{V}^{r}_{t},\forall v\in\mathcal{V}_{exist}^{t}$
Then, since the subgraph nodes $\mathcal{V}^{r}_{t}$ contains both keyword nodes and label nodes, we take the label nodes ${\mathcal{Y}}^{r}_{t}\in\mathcal{V}^{r}_{t}$ in the gernerated sub-graph as our target candidate nodes for the text $t$.
Since this problem is the NP-hard (Su et al., 2020), it is infeasible to obtain the optimal result in polynomial time. Therefore, to solve this problem, we propose a greedy algorithm based on the Melhorn’s algorithm (Mehlhorn, 1988a). Our algorithm generates the candidate label set ${\mathcal{Y}}^{r}_{t}\subseteq\mathcal{Y}^{r}$ for text $t$ at the $r$-th round. The pseudo code of our algorithm is shown in Appendix C, following the Melhorn’s algorithm, our algorithm can also achieve 2-approximate.
Firstly, we calculate the shortest path between each keyword node $v$ in $\mathcal{V}_{exist}^{t}$ to all other nodes in $\mathcal{G}^{r}$ by calculating the minimum spanning tree $MST$ of $\mathcal{G}^{r}$ Here, the keyword nodes $v$ in $\mathcal{V}_{exist}^{t}$ are served as terminal nodes that determines the final generated candidate labels w.r.t. to the weighted graph $\mathcal{G}^{r}$ Secondly, we create a new auxiliary graph $H$ where the edges represent the shortest paths between the closest terminal nodes. Thirdly, we construct the minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$ , and then add the shortest paths between each two nodes in $MST^{\prime}$ to the Steiner Tree $ST$. Lastly our candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ for text $t$ are calculated as the interselect of $ST$ and all target labels $\mathcal{Y}^{r}$ at round $r$: $ST\cap\mathcal{Y}^{r}$. Please refer to Appendix C for more details.
3.4. Part 3: Text Classification and Online Index
In this subsection, we introduce GORAG’s classification part, where it futher enrich the weighted graph by adding query text information with the online indexing mechanism, addressing the narrow source issue. To begin with, we introduce how GORAG performs text classification based the LLM. Specifically, given each unlabled query text $t$, GORAG utilizes LLM to predict its class label $y^{*}$ by constructing an input prompt $c_{t}$ as follows:
(11) $c_{t}=Concat(t,\mathcal{K}^{t},\hat{\mathcal{Y}}^{r}_{t},\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}).$
$c_{t}$ can be considered as the text concatenation of the extracted keywords $\mathcal{K}^{t}$ from the text $t$, and the candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ obtained by Algorithm 2. Also, the $\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}=\{\mathcal{K}_{y_{i}}\}_{i=1}^{|\hat{\mathcal{Y}}^{r}_{t}|}$ and $\mathcal{K}_{y_{i}}$ are the representative keywords of each label $y_{i}\in\hat{\mathcal{Y}}^{r}_{t}$.
(12) $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r}),$
which were generated by LLM with the label’s semantic label name if avaiable. Next, with a classification instruction prompt $p_{classify}$.
(13) $\vspace{-5px}y^{*}_{t}=LLM(p_{classify},c_{t}),$
the LLM would try to select the best-suited label $y^{*}_{t}\in\mathcal{Y}^{t}$ to annotate the text $t$.
To fully leverage the text-extracted keywords, GORAG utilizes an online indexing mechanism to incrementally update keywords that do not yet exist in the weighted graph $\mathcal{G}^{r}$ at the $r$-th round to $\mathcal{G}^{r}$ based on the text $t$’s predicted label $y^{*}_{t}$. To be specific, each keyword node $v\in\mathcal{V}_{notexist}^{t}$ would be added to the original graph’s node set $\mathcal{V}^{r}$ and be assigned with an edge $e_{v,y}$ connecting it with the text $t$’s predicted label $y_{t}^{*}$:
(14) $\vspace{-5px}\mathcal{V}^{r}=\mathcal{V}^{r}\cup\mathcal{V}_{notexist}^{t},\ \mathcal{E}^{r}=\mathcal{E}^{r}\cup\mathcal{E}_{oi}^{t},$
where $\mathcal{E}_{oi}^{t}=\{e_{v,y_{t}^{*}}|v\in\mathcal{V}_{notexist}^{t}\}$ denotes the set of all newly assigned edges $e_{v,y}$ between keyword node $v$ and its predicted label $y_{t}^{*}$, for these newly assigned edges, their weight is calculated with the edge weighting mechanism illustrated in Equation (5).
This online indexing mechanism has two purposes. Firstly, it incrementally enriches the weighted graph $\mathcal{G}^{r}$ by incorporating new, previously unseen keywords, thus expanding the graph’s vocabulary and improving its adaptability. Secondly, by linking these new keywords to their predicted labels with weighted edges, this mechanism captures their relevance and context more effectively, enabling better utilization of the graph for future predictions.
4. Experiments
In this section, we present the experimental evaluation of our framework GORAG on the DFSTC task. We compare the performance of GORAG against six effective baselines spanning three technical categories, using three datasets with distinct characteristics. Specifically, we first outline the experimental setup in Section 4.1, including details on datasets, and evaluation metrics. Due to the space limitation, we put the details of baselines, and hyperparameter configurations in Appendix Section E. Next, we report the experimental results, focusing on both effectiveness and efficiency evaluations, in Section 4.2. Finally, we conduct an ablation study and a case study, presented in Section 4.5 and Appendix Section E.5.
4.1. Experiment Settings
4.1.1. Few-shot setting
In this paper, we employ 1-shot, 5-shot, and 10-shot settings for few-shot training, where each setting corresponds to using 1, 5, and 10 labeled training samples per class, respectively. We omit experiments with more than 10 labeled samples per class, as for WOS dataset, there are already over 1300 labeled training data under 10-shot setting.
4.1.2. Datasets
We select the Web-Of-Science (WOS) dataset (Kowsari et al., 2017) and the Reuters few-shot text classification dataset by (Bao et al., 2020) for evaluating GORAG’s performance. We chose these datasets because they have a considerable amount of classes and text tokens, which is complex and can reflect the real-world scenario. As shown in Table 1, texts from these datasets can contain multiple sentences and have average tokens ranging from 105 to 200 per text. The IFS-Rel dataset was extended from its original version in (Xia et al., 2021), which only has 32 labels in total and an average of 25 tokens per text.
Table 2. Classification accuracy on WOS, Reuters, and IFS-Rel dataset.
Dataset Category Model Round
R1 R2 R3 R4
1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
WOS NN-based Entailment (Xia et al., 2021) 0.3695 0.3823 0.4187 0.3994 0.4471 0.4222 0.4510 0.4857 0.4787 0.4030 0.4387 0.4442
Long Context RAG NaiveRAG (Gao et al., 2023) 0.3885 0.3904 0.3897 0.2267 0.2154 0.2187 0.1821 0.1475 0.1799 0.1653 0.1556 0.1649
Propositionizer (Chen et al., 2023) 0.1241 0.1306 0.1297 0.1074 0.1421 0.1303 0.1771 0.1645 0.1603 0.1611 0.1637 0.1656
Compression RAG LongLLMLingua (Jiang et al., 2024a) 0.3806 0.3823 0.3901 0.2155 0.2202 0.2198 0.1770 0.1567 0.1608 0.1468 0.1382 0.1493
GraphRAG (Edge et al., 2024) 0.3852 0.3897 0.3906 0.2213 0.2197 0.2219 0.1816 0.1770 0.1786 0.1641 0.1634 0.1625
LightRAG (Guo et al., 2024) 0.3930 0.3806 0.3815 0.2202 0.2216 0.2145 0.1743 0.1767 0.1799 0.1625 0.1626 0.1632
GORAG 0.4866 0.4990 0.5134 0.4790 0.5109 0.5284 0.4736 0.4996 0.5234 0.4380 0.4717 0.4929
Reuters NN-based Entailment (Xia et al., 2021) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Long Context RAG NaiveRAG (Gao et al., 2023) 0.0000 0.0000 0.0375 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Propositionizer (Chen et al., 2023) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Compression-RAG LongLLMLingua (Jiang et al., 2024a) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0258 0.0065 0.0085
GraphRAG (Edge et al., 2024) 0.1375 0.1625 0.1000 0.0688 0.0813 0.0500 0.0375 0.0417 0.0417 0.0291 0.0375 0.0375
LightRAG (Guo et al., 2024) 0.0500 0.1375 0.1125 0.0250 0.0813 0.0563 0.0333 0.0333 0.0333 0.0125 0.0333 0.0208
GORAG 0.0875 0.0875 0.1000 0.1750 0.1688 0.1438 0.1667 0.1958 0.2167 0.1667 0.2516 0.2193
IFS-Rel NN-based Entailment (Xia et al., 2021) 0.3391 0.6046 0.5859 0.4008 0.5516 0.5828 0.3061 0.4703 0.5193 0.2971 0.4395 0.4895
Long Context RAG NaiveRAG (Gao et al., 2023) 0.4844 0.4719 0.4828 0.3711 0.3555 0.3656 0.2708 0.2750 0.2625 0.2516 0.2465 0.2527
Propositionizer (Chen et al., 2023) 0.1719 0.1828 0.1938 0.1602 0.1727 0.1750 0.1792 0.1552 0.1770 0.1574 0.1688 0.1664
Compression-RAG LongLLMLingua (Jiang et al., 2024a) 0.4703 0.5203 0.4297 0.3458 0.3055 0.3733 0.2452 0.2206 0.2713 0.2154 0.2029 0.1997
GraphRAG (Edge et al., 2024) 0.4797 0.4656 0.4734 0.3672 0.3641 0.3641 0.2740 0.2776 0.2708 0.2453 0.2465 0.2426
LightRAG (Guo et al., 2024) 0.4484 0.4234 0.4656 0.3531 0.3477 0.3523 0.2760 0.2760 0.2682 0.2340 0.2367 0.2391
GORAG 0.4875 0.5266 0.5313 0.4195 0.3781 0.4374 0.2953 0.2974 0.3328 0.2973 0.2571 0.3034
Table 3. The size of the constructed weighted graph in each round.
Dataset Model Round
R1 R2 R3 R4 After R4
Node # Edge # Node # Edge # Node # Edge # Node # Edge # Node # Edge #
WOS GORAG offline 1,149 1,405 3,681 4,587 4,947 6,204 5,438 6,879 5,438 6,879
GORAG 1,149 1,405 12,108 12,245 25,357 26,290 34,521 35,919 44,283 45,973
Reuters GORAG offline 117 119 192 200 263 284 331 366 331 366
GORAG 117 119 584 594 1,013 1,046 1,342 1,405 1,533 1,600
IFS-Rel GORAG offline 189 213 242 275 372 469 454 594 454 594
GORAG 189 213 5,299 5,503 9,417 9,934 13,293 14,074 16,705 17,486
We split the WOS dataset into 4, 6 and 8 rounds, the Reuters and IFS-Rel datasets into 4 rounds, in order to maintain a comparable number of new label and testing data in each round. Due to the space limitation, we mainly display our experinents on the 4 rounds spit version of both datasets in Table 2, and the experiment result of 6 and 8 rounds are shown in Figure 5(a) and 5(b). Different from the WOS and IFS-Rel datasets, the labels of Reuters are in numeric formats, and their text-formed label names are not available. According to Table 2, this feature makes many baselines fail to carry out classification task on the Reuters dataset. Further statistics of these two dataset being splitted into 4 rounds are shown in Table 1.
4.1.3. Evaluation Metrics
In this paper, aligning with previous DFSTC models (Xia et al., 2021; Meng et al., 2020), we use classification accuracy as the main evaluation metric. Given that LLMs can generate arbitrary outputs that may not precisely match the provided labels, we consider a classification correct only if the LLM’s output exactly matches the ground-truth label name or label number, if the model cannot give out a class within the provided target label set, we denote this classification as a hallucination. We further provide the macro recall score and the hallucination rate on the most complicated and imbalanced test dataset WOS in Table 8 in Appendix.
4.1.4. Baselines
As shown in Table 2, we compare GORAG’s performance with 7 baselines from 3 technical categories for few-shot experiments, they are NN-based Entailment (Xia et al., 2021), Long Context-based NaiveRAG (Gao et al., 2023), Propositionizer (Chen et al., 2023), and Compression-based LongLLMLingua (Jiang et al., 2024a), GraphRAG (Edge et al., 2024), and LightRAG (Guo et al., 2024). Further illustrations of these chosen baselines are listed in subsection E.1 in Appendix. For GraphRAG and NaiveRAG, we use the implementation from (nan, [n.d.]) which optimizes their original code and achieves better time efficiency while not affect the performance; For all other baselines, we use their official implementations.
Also, long text inputs affect LLM performance and increase costs (Chen et al., 2023; Jiang et al., 2024a; Liu et al., 2024a). WOS and Reuters datasets, with lengthy texts and many labels, caused GPU out-of-memory (e.g., llama3-8B, 1-shot, A100-40GB). Thus, we compare open-source LLMs with GORAG in the 0-shot setting (Table 4).
4.2. Few-shot Experiments
4.2.1. Effectiveness Evaluation
As shown in Table 2, GORAG achieves the best classification accuracy over all four rounds with 1-shot and 5-shot labeled indexing data on WOS dataset, it surpasses all RAG-based baselines as well as the state-of-the-art model Entailment. Compared with Entailment, when apply 1-shot setting, GORAG achieves at most 31.6% accuracy gain at the first round. On Reuters dataset, GORAG achieves the best classification accuracy over the last 3 rounds with 1-shot indexing data.
Furthermore, based on the experiment results in Table 2, we can have following observations:
  • Firstly, NN-based Entailment fails on Reuters due to its need for text-formed label names. As noted in prior studies (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023), it overfits easily, with recall and hallucination issues shown in Table 8. Further analysis is in Appendix subsection E.3.
  • Secondly NaiveRAG performs better than Entailment in round 1 of the WOS and IFS-Rel datasets by avoiding overfitting, but its long LLM inputs hinder its ability to adapt to dynamic label updates, leading to a significant performance drop from rounds 2 to 4. While Propositionizer retrieves more fine-grained results, its lack of task awareness results in consistently poor performance on DFSTC tasks. On the Reuters dataset, both Long Context RAG baselines fail when label names are unavailable, as the absence of semantic-rich labels prevents effective classification, especially with long LLM inputs.
  • Thirdly, compression-based RAG models shorten LLM inputs and emphasize labels and key terms but only slightly improve performance across datasets. They struggle with dynamic label updates due to issues like uniform indexing, threshold dependency, and narrow sources, leading to inaccurate retrievals. Additionally, failing to compress label sets results in lengthy LLM inputs as new labels are introduced in each round.
  • Lastly, GORAG addresses the issues of Compression-based RAG models. From R1 to R4, GORAG adapts better to dynamic label updates. On the Reuters dataset, when label names are not available, GORAG achieves comparable or better performance than other baselines. According to Table 3, a vast amount of knowledge is added to the weighted graph by the online indexing mechanism, and contributes to our effectiveness.
4.2.2. Efficiency Evaluation
(a) 1-shot setting.
(b) 5-shot setting.
(c) Time cost on WOS dataset.
(d) Time cost on Reuters dataset.
Figure 5. Experiment result of WOS dataset under 1-shot and 5-shot settings for at most 8 rounds, and model Training/Indexing time cost on two datasets under 1-shot setting.
To evaluate the efficiency of GORAG, we compare GORAG’s indexing time cost with other baselines’ training or indexing time cost on the 4 round split version of WOS and Reuters datasets, the results are shown in Figure 5(c) and 5(d). For both datasets, Propositionizer achieves the worst indexing efficiency. This is because it employs an additional fine-tuned LLM to convert each side information into propositions. In this process, this extra LLM rewrite each labeled or query text and generates a long list of propositions, introducing more efficiency overhead compared to other models that use the LLM solely for classification or keyword extraction, which typically generate only a few tokens. It is also worth noting that, although GORAG’s online indexing mechanism introduces a vast amount of knowledge into the weighted graph, we can still achieve comparable or better time efficiency with other baselines.
4.3. Zero-shot Experiments
To further study GORAG’s 0-shot ability, we compare GORAG’s performance with some widely applied open-source LLM models without RAG approaches. We conducted experiments on the WOS dataset by providing only the label names, without any labeled data to GORAG for creating the graph, the result is shown in Table 4. In the 0-shot setting, GORAG first generates label descriptions based on the label names, then extracts keywords from these descriptions, without using any information from labeled texts. Note that the Entailment model cannot be applied in this setting, as they require labeled data for training. For other RAG-based baselines, they are considered equivalent to their backbone LLMs in the 0-shot setting due to a lack of external information sources.
As shown by the results, GORAG achieves comparable performance with other open-source LLM models in the first round and outperforms them in all subsequent rounds. This is because, when the number of target labels is small, the benefit of compressing the target label set is not as pronounced. However, as the number of labels increases, the importance of filtering the label set becomes more significant, allowing GORAG to consistently outperform other open-source LLM models.
Table 4. Zero-shot experiment result on WOS dataset.
Model Round
R1 R2 R3 R4
Qwen2-7B (Yang et al., 2024) 0.3322 0.1790 0.1451 0.1444
Mistral0.3-7B (Jiang et al., 2023) 0.1436 0.0540 0.0336 0.0270
LLaMA3-8B (Touvron et al., 2023) 0.3351 0.1614 0.1161 0.0930
GORAG 0.3305 0.2567 0.2230 0.2102
4.4. Generalization Evaluation
To evaluate the generalization ability of GORAG, we conducted additional experiments on the 6 and 8 round split versions of the WOS dataset under 1-shot and 5-shot settings, as the WOS dataset contains more testing data and labels than Reuters dataset, and more complex texts than the IFS-Rel dataset. In Figure 5, the results for rounds 1 to 4 are obtained from the 4-round split version, rounds 5 and 6 are obtained from the 6-round split version, and rounds 7 and 8 are obtained from the 8-round split version of the WOS dataset. This setting can ensure a sufficient amount of testing data at each round, As illustrated in Figure 5, the classification accuracy of GraphRAG and NaiveRAG drops significantly from round 1 to 8, while GORAG maintains competitive classification accuracy as the number of rounds increases through 1 to 8, with both 1 and 5-shot setting, demonstrating the importance of a more precise retrieval.
Table 5. Ablation experiments on WOS for GORAG’s variants.
Dataset Model Round
R1 R2 R3 R4
WOS GORAG offline 0.3063 0.2302 0.2455 0.2156
GORAG unit 0.4706 0.4394 0.4407 0.3899
GORAG keyword 0.4746 0.4606 0.4455 0.4030
GORAG 0.4862 0.4649 0.4814 0.4210
Reuters GORAG offline 0.0000 0.1000 0.0917 0.1032
GORAG unit 0.0000 0.0625 0.1667 0.0795
GORAG keyword 0.0000 0.1000 0.1167 0.1438
GORAG 0.0875 0.1750 0.1667 0.1667
4.5. Ablation Study
To analyze GORAG’s superior performance, we conduct an extensive ablation study in the 1-shot setting, as RAG models show minimal performance changes across shot settings. Specifically, we examine how different GORAG components impact classification performance using its variants. GORAG offline removes the online indexing mechanism. GORAG unit removes the edge weighting mechanism, every edge in this variant is assigned with weight 1. GORAG keyword only uses the keyword extracted from the text to create LLM classification input, rather than the whole text.
As shown in Table 5, firstly, GORAG offline achieves the worst accuracy, this is because it lacks a comprehensive retrieval source, making its generated candidate labels inaccurate, hence limits its performance. Secondly, GORAG unit outperforms GORAG offline, but still sub-optimal to GORAG, demonstrates the importance of modeling the varying importance between keywords and each label, which leads to accurate retrievals. Lastly, GORAG keywords demonstrates the important information for text classification are mostly text keywords, as it achieves comparable or better performance than other ablated baselines.
5. Conclusion
In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Text Classification (DFSTC) task. Extensive experiments on three datasets with different characteristics demonstrate the effectiveness of GORAG in classifying texts with only a limited number or even no labeled data. Additionally, GORAG shows its effectiveness in adapting to the dynamic updates of target labels by adaptively retrieving candidate labels to filter the large target label set in each update round. With extensive of ablation studies, we confirm that GORAG’s edge weighting, adaptive retrieval, and online indexing mechanisms contribute to its effectiveness. For future work, we aim to further enhance GORAG’s performance and explore its application in more general scenarios.
Appendix A Notations
Due to the space limitation, we put the summary table on important notations here.
Table 6. Summary on Important Notations.
Notations Meanings
$t$ Text.
$w$ The edge weight.
$r$ The round of label updation.
$y\in\mathcal{Y}^{r}$ The target label at round $r$.
$k$ The number of labeled data provided per target label.
$f^{r}_{\theta}$ The function that learn all target labels’ score for texts.
$\mathbf{s}^{r}$ The score for all target labels at round $r$.
$y^{r*}$ The predicted label for the text at round $r$.
$\mathcal{G}^{r}(\mathcal{V},\mathcal{E})$ weighted graph with node set $\mathcal{V}$ and edge set $\mathcal{E}.$
$\mathcal{Y}^{r}$, $\mathcal{Y}^{r}_{new}$ All target label and new target labels at round $r$.
$\mathcal{D}^{r}$, $\mathcal{D}^{r}_{new}$ The labeled training text for all seen labels at round $r$.
$\mathcal{T}^{r}$, $\mathcal{T}_{new}^{r}$ All seen texts at round $r$.
$\mathcal{T}_{test}^{r}$ The query text for at round $r$.
$\mathcal{K}_{new}^{r}$ The extracted keyword set for new labels of round $r$.
$\mathcal{V}_{new}^{r}$ The set of new graph nodes to be added in round $r$.
$\mathcal{N}(\cdot)$ The neighbor set of a node in the weighted graph.
$p_{extract}$ The extraction instruction prompt.
$p_{gen}$ The generation instruction prompt.
$p_{classify}$ The classification instruction prompt.
$CS(\cdot)$ The correlation score between keywords and labels.
$w^{r}_{v,y}$ The weight of edge $e_{v,y}$ in round $r$.
$P_{u,v}$ The shortest path between node $u$ and node $v$.
$\mathcal{V}_{exist}$ The keywords that exist in graph.
$\mathcal{V}_{notexist}$ The keywords do not exist in graph.
$\mathcal{E}_{oi}^{t}$ Edges added based on text $t$.
Appendix B Complexity Analysis
In this section, we analysis the time and space complexity of GORAG’s graph construction, retrieval and prediction procedure. We denote the maximum number of terms of the input text as $m_{t}$, the number of unique terms of training corpus $\mathcal{D}_{new}^{r}$ as $u$, the LLM’s maximum input token length as $m_{i}$, LLM’s its maximum extraction, generation, and classification token length as $m_{e}$, $m_{g}$, and $m_{c}$, respectively.
B.1. Graph Construction Complexity
Firstly, we analysis the time and space complexity of GORAG’s graph construction. For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e}))$; If label names are available, the time complexity of generat label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}))$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$ times, and merging graph requires
$O(|\mathcal{Y}^{r}_{new}||\mathcal{Y}^{r-1}|)$. Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
$O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e})+|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}+|\mathcal{Y}^{r-1}|)+|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$.
For the space complexity, the graph is stored with the weighted adjency matrix, hence needs $O(|\mathcal{E}^{r}|)$ space; Storing the training corpus at the $r$-th round would need $O(u|\mathcal{D}_{new}^{r}|)$ space; Storing the representive keywords $\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}$ would need $O(|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$ space. Hence the total space complexity of GORAG’s graph construction at the $r$-th round would be:
$O(|\mathcal{E}^{r}|+u|\mathcal{D}_{new}^{r}|+|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$.
B.2. Retrieval and Classification Complexity
The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (Mehlhorn, 1988b); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)$; The time complexity of the final classification by LLM is $O(m_{i}+m_{c})$. Hence, the total time complexity of GORAG’s adaptive retrieval and classification can be denoted as
$O(|\mathcal{T}_{test}^{r}|(|\mathcal{E}^{r}|+|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|+m_{i}+m_{c})).$
For the space complexity, $O(|\mathcal{E}^{r}|)$ to store the graph, and $O(u|\mathcal{T}_{test}^{r}|)$ space is needed to store the testing corpus. Hence, the total space complexity of GORAG’s Retrieval and Classification procedure is
$O(|\mathcal{E}^{r}|+u|\mathcal{T}_{test}^{r}|).$
Appendix C Pseudo Code of GORAG
In this section, we provide the pseudo code of GORAG’s graph construction and adaptive candidate type generation algorithm in 1 and 2, respectively.
Appendix D Hardness Proof of The Adaptive Candidate Type Generation problem
Theorem D.1.
The Adaptive Candidate Type Generation problem is NP-hard.
Proof.
To demonstrate that the Adaptive Candidate Type Generation problem is NP-hard, we provide a simple reduction of our problem from the Steiner Tree problem. Since the Steiner Tree problem is already proven to be NP-hard (Su et al., 2020), we show that there is a solution for the Steiner Tree problem if and only if there is a solution for our problem. Firstly, given a solution $S$ for our problem, we can construct a Steiner Tree by generating a minimum spanning tree for all nodes in $S$ on graph $\mathcal{G}^{r}$ then connecting all nodes from $\mathcal{V}_{exist}^{t}$ to their closest neighbor nodes in $S$. Secondly, any Steiner Tree $ST$ that contains any node $y\in\mathcal{V}^{r}_{t}$ is also a solution of our problem with |$|\{y\in\mathcal{V}^{r}_{t}\}|$ nodes. Thus, we prove that the adaptive candidate type generation problem is NP-hard. ∎
Algorithm 1 Graph Construction Algorithm of GORAG at round $r$
0:  $\mathcal{D}_{new}^{r}$: Training text set at the $r$-th round and the provided label $y_{i}\in\mathcal{Y}_{new}^{r}$ for each training text $t_{j}$; $p_{extraction},\ p_{gen}$: The extraction and description generation instruction prompt; $\mathcal{G}^{r-1}$: Weighted graph of the previous round $r-1$.
0:  $\mathcal{G}^{r}$: Constructed weighted graph at the $r$-th round.
1:  Let $\mathcal{V}_{text}^{r}=\emptyset$, $\mathcal{V}_{label}^{r}=\emptyset$, $\mathcal{E}_{new}^{r}=\emptyset$
2:  for each text $t\in\mathcal{T}_{new}^{r}$ do
3:     $\mathcal{V}_{text}^{r}=\mathcal{V}_{text}^{r}\cup LLM(p_{extract},t)\cup y_{t}$
4:  end for
5:  if label names are available then
6:     for each label $y_{i}\in\mathcal{Y}_{new}^{r}$ do
7:        Get label description $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r})$.
8:        $\mathcal{V}_{label}^{r}=LLM(p_{extract},\mathcal{K}_{y_{i}})\cup\mathcal{V}_{label}^{r}$
9:     end for
10:  end if
11:  Let $\mathcal{V}_{new}^{r}=\mathcal{V}_{text}^{r}\cup\mathcal{Y}^{r}_{new}\cup\mathcal{V}_{label}^{r}$ be the new node set at the $r$-th round.
12:  for each node $v\in\mathcal{V}_{new}^{r}$ do
13:     $\mathcal{E}_{new}^{r}=\mathcal{E}_{new}^{r}\cup e_{v,y_{i}}$
14:     $\mathcal{W}_{new}^{r}=\mathcal{W}_{new}^{r}\cup w_{v,y_{i}}^{r}$
15:  end for
16:  Let $\mathcal{G}^{r}_{new}=(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r})$ be the newly constructed graph for at the $r$-th round.
17:  Let $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1})$ be the final constructed graph at the $r$-th round.
18:  return  $\mathcal{G}^{r}$
Appendix E Experiment Details
E.1. Further introduction of chosen baselines
NN-based Models
  • Entailment (Xia et al., 2021): Entailment concatenates the text data with each of the label names to form multiple entailment pairs with one text sample, hence increasing the number of training data and enhance its finetuning of a RoBERTa PLM (Liu, 2019). Entailment then carry out classification based on these entailment pairs and convert the text classification task into a binary classification task, for detecting whether the entailment pair formed is correct. When all entailment pairs of a particular text are considered to be incorrect, Entailment will directly give up the classification of this text. In this paper, we consider this classification as a hallucination.
Long Context RAG Models
  • NaiveRAG (Gao et al., 2023): NaiveRAG acts as a foundational baseline of current RAG models. When indexing, it stores text segments of the labeled texts in a vector database using text embeddings. When querying, NaiveRAG generates query texts’ vectorized representations to retrieve side information based on the highest similarity in their embeddings.
  • Propositionizer (Chen et al., 2023): Propositionizer applies a fine-tuned LLM to convert side information into atomic expressions, namely propositions, to facilitate more fine-grained information retrieval than NaiveRAG.
Compression RAG Models
  • LongLLMLingua (Jiang et al., 2024a): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
  • GraphRAG (Edge et al., 2024): GraphRAG is a graph-Based RAG model that employs the LLM to extract texts’ entities and relations, which are then represented as nodes and edges in the weighted graph. GraphRAG then aggregates nodes into communities, and generates a comprehensive community report to encapsulate global information from texts.
  • LightRAG (Guo et al., 2024): LightRAG skips the GraphRAG’s formulation of graph communities and directly retrieve nodes or pathes from created graphs. It calculates the embedding similarity of text extracted entities with graph nodes, to achieve a one-to-one mapping from keywords to graph nodes and retrieve all triples involved these nodes.
Algorithm 2 Adaptive Candidate Type Generation Algorithm
0:  $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$: The constructed weighted graph; $\mathcal{V}_{exist}^{t}\subseteq\mathcal{V}^{r}$: A set of keyword nodes extracted from text $t$ and can be mapped to graph $\mathcal{G}$; $\mathcal{Y}^{r}$: The target label set at the $r$-th round.
0:  $\hat{\mathcal{Y}}^{r}_{t}$: The candidate type retrieved for text $t$.
1:  Compute a minimum spanning tree $MST$ of the graph $\mathcal{G}$.
2:  Let $\mathcal{V}_{term}=\mathcal{V}_{exist}^{t}\cap\mathcal{V}(MST)$ be the terminal nodes in $MST$
3:  Construct a weighted auxiliary graph $H$: $\mathcal{V}(H)=\mathcal{V}_{term}$
4:  for each pair of terminals $u,v\in\mathcal{V}_{term}$ do
5:     Find the shortest path $P_{uv}$ in $MST$ from $u$ to $v$
6:     Let $w^{H}_{u,v}=\sum_{e\in P_{uv}}w_{e}$
7:  end for
8:  Compute a minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$
9:  Let $ST=\emptyset$
10:  for each edge $e_{u,v}\in MST^{\prime}$ do
11:     Add the shortest path $P_{uv}$ in $MST$ to $ST$
12:  end for
13:  Let $\hat{\mathcal{Y}}^{r}_{t}=ST\cap\mathcal{Y}^{r}$
14:  return  $\hat{\mathcal{Y}}^{r}_{t}$
E.2. Hyperparameter and Hardware Settings
In each round of our experiments, we handle NN and RAG-based models differently for training and indexing. For NN-based models, we train them using all labeled data from the current round and all previous rounds. For RAG-based models, we index the labeled data of the current round to the information source from the previous round. The RAG-based model is initialized from scratch only in the first round.
After training or indexing in each round, we test the models using the testing data from the current round and all previous rounds. For GORAG, which employs an online indexing mechanism, we first test its performance with online indexing on the current round’s testing data. We then test its performance on the testing data from all previous rounds without online indexing. This approach allows us to study the effect of online indexing on the performance of later round data when applied to previous round data.
For the hyperparameter settings, we train the Entailment model with the RoBERTa-large PLM for 5 epochs with a batch size of 16, a learning rate of $1\times 10^{-6}$, and the Adam optimizer (Kingma, 2014), following the exact setting in Entailment’s original paper. For LongLLMLingua, we test the compression rate within ${0.75,0.8,0.85}$ and select 0.8, as it achieves the best overall classification accuracy. For GraphRAG and LightRAG, we use their local search mode, as it achieves the highest classification accuracy on both the WOS and Reuters datasets. For GORAG and all RAG-based baselines, we use LLaMA3 (Touvron et al., 2023) as the LLM backbone unless otherwise specified.
All experiments are conducted on an Intel(R) Xeon(R) Gold 5220R @ 2.20GHz CPU and a single NVIDIA A100-SXM4-40GB GPU.
E.3. Further Analysis based on Precision and F1
According to Table 8 and Table 2, although Entailment achieves the second best accuracy than GORAG on WOS dataset, it has a high hallucination rate. This is because the ground truth label of Entailment’s entailment pairs are greatly unbalanced, hence make the model easy to overfit, and tend to predict all entailment pairs of a particular text as incorrect under few-shot settings.
(a) GORAG’s candidate selection algorithm returns candidate labels adaptively based on keywords extracted from the tesing texts.
(b) GORAG’s candidate selection algorithm only returns the ground truth label of the query text.
Figure 6. Two case of GORAG successfully generate candidate labels to help text classification on WOS dataset.
E.4. GORAG with different backbone LLMs
Table 7. Evaluation on different LLM backbones.
Dataset Model Round
R1 R2 R3 R4
WOS GORAG LLaMA3 0.4862 0.4649 0.4814 0.4210
GORAG Qwen2.5 0.5101 0.4839 0.4823 0.4235
In Table 7, we present the results of GORAG with different backbone LLMs, we test GORAG with more advanced Qwen2.5-7B (Qwen Team, 2024), the result demonstates the GORAG’s performance can be further improved with more advanced LLM backbones.
E.5. Case Study
In this section, to illustrate the strength of GORAG’s adaptive retrieval and candidate label generation procedure, we dig into two testing cases select from WOS dataset where GORAG’s adaptive retrieval helps to reduce the LLM input length and benefit text classification. As shown in Figure 6(a), for the case whose ground truth label is Algorithm Design, the candidate label retrieved by GORAG are Computer programming and Algorithm design. As a result, GORAG’s adaptive retrieval algorithm successfully filter the target label set to only contains two candidate labels, and successfully cut down the target label number from at most over 100 to only 2 candidate labels that it consided as the possible labels for the text, significantlly reduce the LLM’s input length and mitigate the lost-in-the-middle issue.
On the other hand, for the case in Figure 6(b) whose ground truth label is Electrical Circuits, since the ground truth label name Electrical Circuits already exists in the text and being extracted as keyword, GORAG’s adaptive retrieval algorithm is almost certain to classify this text into class Electrical Circuits, hence GORAG’s adaptive retrieval and candidate type generation algorithm would only select the only candidate label retrieved from the constructed graph and input to the LLM.
Table 8. The macro recall score and hallucination rate (Hall. Rate) on WOS dataset.
Metrics Category Model Round
R1 R2 R3 R4
1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
Recall NN-based Entailment 0.3502 0.3475 0.3908 0.3021 0.3415 0.3442 0.2327 0.2406 0.2432 0.1789 0.1792 0.1837
Long Context RAG NaiveRAG 0.3324 0.3417 0.3398 0.1615 0.1652 0.1650 0.0943 0.1013 0.0872 0.0568 0.0805 0.0628
Propositionizer 0.1032 0.1131 0.1114 0.0947 0.1210 0.1187 0.1352 0.1298 0.1254 0.1373 0.1345 0.1402
Compression RAG LongLLMLingua 0.3139 0.3105 0.3258 0.1349 0.1388 0.1350 0.0862 0.0837 0.0872 0.0616 0.0610 0.0628
GraphRAG 0.3791 0.3646 0.3732 0.1535 0.1519 0.1557 0.1009 0.1014 0.1015 0.0902 0.0940 0.0912
LightRAG 0.3462 0.3441 0.3519 0.1544 0.1507 0.1466 0.1241 0.1257 0.1226 0.1129 0.1180 0.1101
GORAG 0.4387 0.4659 0.4874 0.3074 0.3472 0.3483 0.2371 0.2513 0.2445 0.1960 0.1970 0.2013
Hal. Rate NN-based Entailment 54.53% 58.96% 51.63% 44.76% 36.62% 37.76% 32.57% 28.14% 27.35% 29.68% 26.86% 24.13%
Long Context RAG NaiveRAG 3.00% 2.32% 2.63% 6.45% 6.16% 6.28% 6.93% 7.46% 7.52% 9.96% 9.50% 9.83%
Propositionizer 6.85% 5.88% 6.97% 7.40% 6.52% 6.09% 5.97% 6.44% 6.30% 5.78% 6.11% 5.52%
Compression-RAG LongLLMLingua 6.50% 6.29% 5.63% 16.45% 16.05% 16.28% 19.63% 20.47% 20.52% 21.96% 21.35% 21.83%
GraphRAG 2.40% 2.23% 2.07% 6.03% 5.97% 6.18% 7.20% 7.58% 7.03% 6.90% 6.39% 6.49%
LightRAG 1.70% 1.24% 1.37% 2.59% 2.78% 3.21% 4.66% 4.55% 4.58% 5.89% 5.80% 5.73%
GORAG 0.66% 1.16% 0.74% 0.30% 0.68% 0.65% 0.41% 0.57% 0.68% 0.40% 0.43% 0.64%
Table 9. The macro recall score and hallucination rate (Hall. Rate) on IFS-Rel dataset.
Metrics Category Model Round
R1 R2 R3 R4
1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
Recall NN-based Entailment 0.3191 0.6015 0.5515 0.2172 0.3043 0.3330 0.1225 0.1802 0.2081 0.1093 0.1267 0.1490
Long Context RAG NaiveRAG 0.4039 0.4194 0.4292 0.2320 0.2009 0.2115 0.1147 0.1163 0.1097 0.0871 0.0895 0.0912
Propositionizer 0.1053 0.1146 0.1104 0.1124 0.1201 0.1197 0.0824 0.1029 0.0902 0.0744 0.0717 0.0799
Compression RAG LongLLMLingua 0.4426 0.4625 0.4044 0.1879 0.1559 0.2158 0.0947 0.0906 0.1113 0.0698 0.0822 0.0696
GraphRAG 0.4264 0.4139 0.4208 0.2116 0.2008 0.2047 0.1213 0.1179 0.1185 0.0906 0.0857 0.0883
LightRAG 0.4221 0.3985 0.4382 0.2344 0.2150 0.2429 0.1270 0.1313 0.1276 0.0902 0.0926 0.0918
GORAG 0.4588 0.4681 0.5000 0.2348 0.2175 0.2405 0.1279 0.1295 0.1290 0.0950 0.0886 0.0919
Hal. Rate NN-based Entailment 8.10% 15.63% 4.69% 26.41% 19.14% 6.25% 26.56% 17.86% 6.67% 6.48% 5.90% 3.28%
Long Context RAG NaiveRAG 1.56% 4.25% 1.78% 3.05% 2.97% 3.13% 3.02% 3.70% 3.54% 4.02% 4.18% 3.63%
Propositionizer 7.82% 7.34% 11.25% 5.78% 6.40% 6.72% 4.84% 6.09% 7.03% 8.13% 8.59% 6.72%
Compression-RAG LongLLMLingua 3.59% 2.97% 6.56% 2.12% 6.23% 5.47% 6.36% 6.00% 3.50% 3.29% 5.65% 8.88%
GraphRAG 4.25% 1.56% 1.25% 2.66% 3.98% 2.89% 3.39% 3.75% 3.80% 3.91% 3.24% 3.09%
LightRAG 3.28% 2.66% 2.03% 2.73% 2.97% 2.27% 1.98% 1.56% 1.67% 2.46% 2.23% 2.30%
GORAG 0.31% 0.63% 0.16% 1.78% 2.81% 2.58% 1.30% 0.78% 1.09% 1.48% 1.02% 1.25%

### v4

@@
-GORAG: Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification
+GORAG: Graph-based Online Retrieval Augmented Generation for Dynamic Few-shot Social Media Text Classification
 Abstract.
-Text classification is a fundamental task in data mining, pivotal to various applications such as tabular understanding and recommendation.
-Although neural network-based models, such as CNN and BERT, have demonstrated remarkable performance in text classification, their effectiveness heavily relies on abundant labeled training data.
-This dependency makes these models less effective in dynamic few-shot text classification, where labeled data is scarce, and new target labels frequently appear based on application needs.
-Recently, large language models (LLMs) have shown promise due to their extensive pretraining and contextual understanding ability.
-Current approaches provide LLMs with text inputs, candidate labels, and additional side information (e.g., descriptions) to classify texts.
-However, their effectiveness is hindered by the increased input size and the noise introduced through side information processing.
-To address these limitations, we propose a graph-based online retrieval-augmented generation framework, namely GORAG, for dynamic few-shot text classification.
-Rather than treating each input independently, GORAG constructs and maintains a weighted graph by extracting side information across all target texts.
-In this graph, text keywords and labels are represented as nodes, with edges indicating the correlations between them.
-To model these correlations, GORAG employs an edge weighting mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a minimum-cost spanning tree tailored for each text input.
-Empirical evaluations demonstrate that GORAG outperforms existing approaches by providing more comprehensive and precise contextual information.
-Large language models, online retrieval-augmented generation, few-shot learning
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/02
-Figure 1. An example of the DFSTC task with two rounds. In round 1, an example text is classified as Machine Learning. In round 2, an example text is classified as Computer Vision.
+Text classification is vital for Web for Good applications like hate speech and misinformation detection.
+However, traditional models (e.g., BERT) often fail in dynamic few-shot settings where labeled data are scarce, and target labels frequently evolve.
+While Large Language Models (LLMs) show promise in few-shot settings, their performance is often hindered by increased input size in dynamic evolving scenarios.
+To address these issues, we propose GORAG, a Graph-based Online Retrieval-Augmented Generation framework for dynamic few-shot text classification.
+GORAG constructs and maintains a weighted graph of keywords and text labels, representing their correlations as edges.
+To model these correlations, GORAG employs an edge weighting mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a tailored minimum-cost spanning tree for each input.
+Empirical evaluations show GORAG outperforms existing approaches by providing more comprehensive and precise contextual information.
+Our code is released at: https://github.com/Wyb0627/GORAG.
+Large Language Model, Retrieval-Augmented Generation, Few-shot Text Classification, Hate Speech Detection, Social Media
+journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3793011isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Social networksccs: Information systems Clustering and classificationccs: Information systems Retrieval models and ranking
 1. Introduction
-Text classification is a fundamental task of text data mining and is connected to various tasks, such as tabular understanding (Chen et al., 2024c; Wang et al., 2024c; Wen et al., 2024a, b; He et al., 2025; Zheng et al., 2023), and recommendation (Du et al., 2024a; Han et al., 2024b; Wang et al., 2024b; Li et al., 2023; Zheng et al., 2024; Chiang et al., 2023; Du et al., 2024b).
-In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Prieto et al., 2016; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters et al., 2018; Wang et al., 2018; Yang et al., 2019; Miao et al., 2021), such as CNN (Kim, 2014), Bert (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on text classification tasks.
-However, the effectiveness of these NN-based approaches depends on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018).
-Furthermore, in real-world applications, such as Web Of Science (Kowsari et al., 2017) and IMDb (imd, [n.d.]), target labels for text often change based on the application’s requirements (Xia et al., 2021), leading to the dynamic few-shot text classification (DFSTC) task.
-As shown in Figure 1, the DFSTC task begins with a few initial classes (e.g., Machine Learning).
-As new topics like Computer Vision emerge in later rounds, models must adapt to these changes and accurately classify new examples with minimal labeled data.
-Hence, how to develop methods to dynamically classify text with limited labeled data available remains a necessity and open problem.
-Depending the technique for the DFSTC task, current models can be mainly categorized into two types, i.e., data augmentation-based models and RAG-based models.
-Firstly, data augmentation-based models (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021) create additional training data by mixing the pairs of the few-shot labeled data and assigning mixed labels indicating the validity for these created data based on the labels of each data pairs (Xia et al., 2021).
-Also, several data augmentation approaches create extra semantic-related content based on the label names (Meng et al., 2018; Meng et al., 2020) to enrich the training corpus further.
-However, due to the limited labeled data, and the label names may not always be available, the generated text data can have very limited patterns.
-Consequently, the text classification models trained on these generated data may over-fit on limited text data and are not generalizable (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023).
-Figure 2.
-An overview of GraphRAG (Edge et al., 2024).
-It first extracts and indexes entities and relations as graph nodes and edges.
-Then it performs Top-K similarity-based retrieval for query-related subgraphs.
-Finally, it classifies texts based on the subgraph summary and the query text.
-Recently, large language models (LLMs) (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), pre-trained on extensive corpora, have achieved significant success due to their superior and comprehensive text understanding abilities.
-However, researchers (Gao et al., 2023; Peng et al., 2024) discover that LLMs can struggle to understand their inputs due to a lack of task-specific knowledge.
-Therefore, the Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024) were proposed to combine the predicted text, candidate labels, and the retrieved side information (e.g., descriptions of text and labels, or documents) as LLM inputs, enable LLMs to understand the predicted text and labels more comprehensively with the help of side information from external sources.
-However, the incorporation of side information can further increase the input size and noise (Chen et al., 2023), which impedes the efficiency and effectiveness of LLMs (Sarthi et al., 2024).
-As a result, Compression-based RAG approaches (Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024) are proposed to compress the length of context and extract the key information.
-Specifically, LongLLMLingua (Xia et al., 2021) and CompaAct (Yoon et al., 2024) propose filtering uninformative texts based on the query.
-However, these approaches may result in information loss and degrade overall performance.
-Then, to address this, graph RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024; Peng et al., 2024; Zhang et al., 2025; Han et al., 2024a), such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2024), propose to to index the unstructured texts into the structured graphs.
-For example, in Figure 2, the basic idea of graph RAG considers text entities as graph nodes, and text sentences describing relations between these entity nodes as graph edges, later, it retrieves subgraphs base on the query text, and generates short summaries of these subgraphs as the side information for LLMs, reducing the input length for LLMs.
-However, existing Compression RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) still have three issues in the DFSTC task.
+Figure 1. An example of DFSTC task with 2 rounds. In Round 1, the text is classified as Threatening Language. In Round 2, another text is classified as a new label: Identity Insults.
+Text content classification is a fundamental task for various social media applications, such as hate speech detection (Geissler et al., 2025; Hee et al., 2024; Chen et al., 2024b; Djuric et al., 2015; Subramanyam et al., 2024; Mandal et al., 2024; Cao et al., 2024; Ji et al., 2024; Guo et al., 2024), misinformation/disinformation detection (Luo et al., 2024; Nesterov et al., 2024; Liu et al., 2025; Toughrai et al., 2025; Cinus et al., 2025; Ashkinaze et al., 2024; Chen et al., 2022; Shang et al., 2024; Ye et al., 2024) , and suicide risk detection (Kang et al., 2024; Li et al., 2025; Naseem et al., 2025; Ge et al., 2025).
+In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters and others, 2018; Wang et al., 2018; Yang et al., 2019), such BERT (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on social media text content classification tasks.
+However, the effectiveness of these NN-based approaches (Johnson and Zhang, 2015; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters and others, 2018; Wang et al., 2018; Yang et al., 2019) depends on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018; Nguyen and Rudra, 2024).
+Furthermore, in real-world applications, such as Abuse Detection (Vidgen et al., 2021), new kinds of social media abuse can frequently appear with the changing needs of user analytics and progress in social research development, introducing the need to classify texts into new, extra labels (Xia et al., 2021), leading to the Dynamic Few-shot Social media Text Classification (DFSTC) task.
+As shown in Figure 1, the DFSTC task starts with an initial set of classes (e.g., Threatening Language).
+As new classes like Identity Insults emerge in later rounds, it becomes challenging to obtain extensive labeled data for these new categories.
+Models must therefore adapt to such changes and accurately classify new examples with limited labeled data.
+Developing methods to dynamically classify text with limited labeled data remains a crucial and open problem.
+Depending on the technique for the DFSTC task, current models can be mainly categorized into two types, i.e., data augmentation-based models and RAG-based models.
+Firstly, data augmentation-based models (Meng et al., 2018, 2020; Xia et al., 2021) create additional data for fine-tuning small language models (e.g., RoBERTa (Liu, 2019)), by mixing the pairs of the few-shot labeled data and assigning mixed labels indicating the validity for these created data based on the labels of each data pairs (Xia et al., 2021).
+However, due to the limited labeled data, and the label names may not always be available; furthermore, the additional data created can have very limited patterns.
+Consequently, the classification models trained with these created data may over-fit and are not generalizable (Liu et al., 2023, 2024a; Lei et al., 2023).
+Recently, Large Language Models (LLMs) (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024a) have excelled in social media text classification through their superior understanding abilities.
+However, they often struggle with task-specific knowledge (Gao et al., 2023; Zhang et al., 2023) and lack accurate reasoning clues to explain responses (Peng et al., 2024; Wang et al., 2024b; Dong et al., 2025; Qian et al., 2025).
+In domains like hate speech or misinformation detection, these flaws can lead to undetected harmful content or wrongful penalties.
+Consequently, Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Xia et al., 2021; Yoon et al., 2024) were proposed to integrate retrieved side information (e.g., label descriptions) as external clues for more accurate classification.
+However, the incorporation of side information can further increase the input size (Chen et al., 2023), which often overwhelms critical details with noises.
+As the key clues of hate speech or misinformation are usually short sentences within long contexts, the ignorance of fine-grained, entity-level information can degrade LLM classification performance.
+Although proprietary LLMs (Hurst et al., 2024) can handle such extended inputs, they inevitably suffer from high token costs.
+To achieve finer-grained RAG, and better preserve the entity-level information, Graph-based RAG approaches (Edge et al., 2024; Guo et al., 2025; Jimenez Gutierrez et al., 2024; Peng et al., 2024; Zhang et al., 2025; Han et al., 2024; gutiérrez2025ragmemory), such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2025), propose to index the unstructured texts into the structured graphs.
+These approaches extract entities and relations from texts as graph nodes and edges, and their retrievals are based on the graph components mapped from the query text.
+However, existing Graph-based RAG approaches (Edge et al., 2024; Guo et al., 2025; Jimenez Gutierrez et al., 2024) still have three issues in the DFSTC task.
   1. (1)
     Uniform-indexing issue: Some of these approaches construct graphs by indexing extracted text chunks uniformly.
-They do not consider the varying importance and extraction confidence of each text chunk, which may provide incorrect and unreliable context for LLMs.
+They do not consider the varying importance of each text and extraction confidence of keyword entity.
+For example, in Figure 1, keyword lazy is related to all round 2 labels with different importance.
+Hence, uniform-indexing may provide inaccurate graph context for LLM classification.
   2. (2)
-    Threshold-dependent issue: These approaches select relevant information for each input text based on a globally predefined threshold.
-However, the optimal retrieval threshold can vary across different text samples; hence, relying on a globally predefined threshold can be suboptimal for the entire dataset.
+    Non-adaptive Retrieval issue: These approaches select relevant information for each input text based on a globally predefined threshold.
+However, the optimal retrieval threshold can vary across different query text and different rounds; hence, this non-adaptive retrieval approach can be suboptimal.
   3. (3)
-    Narrow-source issue: These approaches only retrieve side information from the each few-shot training text and ignore the important information among different the query texts, consequently limiting their retrieve accuracy and comprehensiveness.
-To address these issues, we propose a novel hyperparameter-free Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot text classification, called GORAG.
+    Narrow-source issue: These approaches only retrieve side information from a limited number of few-shot training text and ignore the important information among query texts.
+Since the new keywords related to certain labels in hate speech or misinformation detection emerge frequently in social media, this issue consequently limits models’ retrieval comprehensiveness.
+To address these issues, we propose a novel Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot social media text classification, called GORAG.
 In general, GORAG constructs and maintains an adaptive online weighted graph by extracting side information from all target texts, and tailoring the graph retrieval for each input.
-Firstly, GORAG constructs a weighted graph with keywords extracted from labeled texts as keyword nodes, and text labels as label nodes.
-The edge of this graph represents the relationship between keywords and labels.
-Secondly, GORAG employs an edge-weighting mechanism to assign different weights to edges created by the last step.
-An example of the constructed weighted graph is shown in Figure 3.
-The edge weights represent the keywords’ importance and relevance to the respective text’s label.
-Thirdly, GORAG constructs a minimum-cost spanning tree on the constructed weighted graph based on keywords of the query text, and then retrieves the label nodes within the spanning tree as candidate labels.
-Since the generated spanning tree is determined solely by the constructed graph and the keywords of the query text, GORAG achieves adaptive retrieval without relying on any human-defined retrieval thresholds, making the retrieval more precise regarding each query text.
-Lastly, GORAG applies an online indexing mechanism, which can enrich the weighted graph with the keywords from query texts, enhancing its few-shot learning performance by providing a more comprehensive retrieve source.
+Firstly, GORAG constructs a weighted graph with keywords extracted from labeled texts as keyword nodes, and text labels as label nodes, to help the model better capture graph topology (Yang et al., 2024b; Zhao et al., 2024; Lin et al., 2024).
+To model relationships, GORAG employs an edge-weighting mechanism to assign different weights to edges, representing keywords’ importance and relevance to the respective text’s label.
+Secondly, GORAG constructs a minimum-cost spanning tree based on query keywords and retrieves the label nodes within it as candidate labels.
+Since this MST is determined solely by the constructed graph and the keywords of the query text, GORAG achieves adaptive retrieval without human-defined thresholds, ensuring precision for each input.
+Finally, GORAG applies an online indexing mechanism to enrich the graph with keywords from query texts.
+By calculating weights that reflect the semantic relevance of keywords to labels, GORAG effectively reduces the potential noise introduced by inaccurate LLM classification during real-time updates, ensuring the graph remains a robust and comprehensive retrieval source
 We summarize our novel contributions as follows.
-  • We present a hyperparameter-free online RAG framework for DFSTC tasks, namely GORAG.
-The GORAG framework consists of three steps: graph construction, candidate label retrieval, and text classification, aim at addressing the uniform-index, threshold-dependent and narrow-source issues.
+  • We present an online RAG framework for DFSTC tasks, namely GORAG, which consists of three steps: Graph Construction, Candidate Label Retrieval, and Text Classification, aiming to address the Uniform-indexing, Non-adaptive Retrieval, and Narrow-source issues.
   • We develop a novel graph edge weighting mechanism based on the text keywords’ importance within the text corpus, which enables our approach to effectively model the relevance between keywords and labels, thereby helping the more accurate retrieval.
-  • To avoid any human-selected thresholds, we formulate the candidate label retrieval problem which is akin to the NP-hard Steiner Tree problem.
-To solve this problem, provide an efficient and effective solution to generate candidate labels.
+  • To achieve adaptive retrieval, we formulate the candidate label retrieval problem which is akin to the NP-hard Steiner Tree problem, and provide an efficient and effective solution based on the greedy algorithm.
+  • Extensive experiments from the social media contextual abuse benchmark and the web news misinformation detection benchmark demonstrate the effectiveness and efficiency of GORAG.
 2. Preliminary and Related Works
-We first introduce the preliminaries of dynamic few-shot text classification in Section 2.1 and then discuss the related works for DFSTC task in Section 2.2.
+We first introduce the preliminaries of Dynamic Few-shot Social media Text Classification (DFSTC) in Section 2.1 and then discuss the related works for DFSTC task in Section 2.2.
 The important notations used in this paper are listed in Table 7 in Appendix A.
-2.1. Dynamic Few-shot Text Classification Task
-Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world application that involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$.
+2.1. DFSTC Task
+Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world web for good applications, including social media hate speech detection (Geissler et al., 2025; Hee et al., 2024; Chen et al., 2024b; Djuric et al., 2015; Subramanyam et al., 2024; Mandal et al., 2024; Cao et al., 2024; Ji et al., 2024; Guo et al., 2024), misinformation/disinformation detection (Luo et al., 2024; Nesterov et al., 2024; Liu et al., 2025; Toughrai et al., 2025; Cinus et al., 2025; Ashkinaze et al., 2024; Chen et al., 2022; Shang et al., 2024; Ye et al., 2024).
+It involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$.
 Traditional approaches require fine-tuning with large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available.
-Recently, few-shot learning addresses this limitation by enabling models to classify text with only a small number of labeled examples per class (Meng et al., 2018; Meng et al., 2020).
-On the other hand, dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024d) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones.
-Combining these aspects, Dynamic Few-Shot Text Classification (DFSTC) task allows the model to handle evolving classification tasks with minimal labeled data, and better suited for real world scnerio (Xia et al., 2021).
+Recently, few-shot learning addresses this limitation by enabling models to classify social media texts into various classes with only a small number of labeled examples per class (Meng et al., 2018, 2020).
+On the other hand, dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024a) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones, as frequently involve extensive human labeling for these new classes can be difficult.
+For example, in Figure 1, new and more fine-grained classes of hate speech, such as Identity Insults, can serve as complements to categories that already exists.
+Combining these aspects, Dynamic Few-Shot Social media Text Classification (DFSTC) task allows the model to handle evolving classification tasks with minimal labeled data, and is better suited for real world scenarios (Xia et al., 2021).
 In DFSTC task, the model is provided with multiple rounds of new class updates.
-Specifically, In each round $r$, a new set of classes $\mathcal{Y}_{new}^{r}$ is introduced, and the labelled dataset for $\mathcal{Y}_{new}^{r}$ is denoted as $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{new}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y_{i}\in\mathcal{Y}_{new}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$.
-Also, we denote candidate cumulative labels from the first round to the $r$-th round as $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}_{new}^{i}$.
-Formally, at the $r$-th round, given the candidate labels $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}_{new}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}={f}^{r}_{\theta}(t_{u},\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{u}$.
+Specifically, in each round $r$, a set of new labels $\mathcal{Y}^{r}_{new}$ is introduced.
+The complete candidate label set for round $r$, denoted as $\mathcal{Y}^{r}$, is the union of all new labels from rounds $i=1$ to $r$: $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}^{i}_{new}$.
+In round $r$, the labeled training set for class labels in $\mathcal{Y}^{r}$ is denoted as $\mathcal{D}^{r}=\cup_{y_{i}\in\mathcal{Y}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y_{i}\in\mathcal{Y}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$.
+Formally, at the $r$-th round, given the labels in $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}={f}^{r}_{\theta}(t_{u},\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{u}$.
 Then, we can get the predicted label $y^{r*}\in\mathcal{Y}^{r}$ for the unseen text $t_{u}$ as follows.
-(1) $y^{r*}=\arg\max_{y^{r}\in\mathcal{Y}^{r}}({f}_{\theta}^{r}(t_{u},\mathcal{Y}^{r})[y^{r}])$
+(1) $y^{r*}=\arg\max_{y^{r}\in\cup_{i=1}^{r}\mathcal{Y}^{i}}({f}_{\theta}^{r}(t_{u},\cup_{i=1}^{r}\mathcal{Y}^{i})[y^{r}])$
 DFSTC is valuable for its ability to learn from limited labeled data, adapt to evolving class distributions, and address real-world scenarios where classes and data evolve over time.
-2.2. Dynamic Few-shot Text Classification Model
-Current DFSTC models can be broadly categorized into NN-based models and RAG-based models, RAG-based models can be further categorized into Long Context RAG models, and Compression-based RAG models.
+2.2. DFSTC Models
+Current Dynamic Few-Shot Social media Text Classification (DFSTC) models can be broadly categorized into NN-based models and RAG-based models, RAG-based models can be further categorized into Long Context RAG, and Graph-based RAG models.
 2.2.1. NN-based models
-NN-based models (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021) generate additional data contrastively based on the few-shot labeled data and the text formed label names to train the NN-based classifier.
-However, due to the limited labeled data, the generated text data of these models can have very limited patterns, which makes them prone to overfitting (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023).
-Also, as the text-formed label names are not always available in real-world scenarios, these NN-based models may not always be applicable.
-Figure 3.
-An example of the constructed weighted graph, where red nodes denote the label nodes and black nodes denote the keyword nodes.
-The lower the edge cost is, the more related the keyword is to the respective label.
+NN-based models (Meng et al., 2018, 2020; Xia et al., 2021) generate additional data contrastively based on the few-shot labeled data and the text formed label names to train the NN-based classifier.
+However, due to the limited labeled data, the generated text data of these models can have very limited patterns, which makes them prone to overfitting (Liu et al., 2023, 2024a; Lei et al., 2023).
+In web for good applications such as hate speech detection, different kinds of hate speech contents are usually a minority compared with neutral contents, this highly unbalanced data distribution can further overfit the NN-based models and affect their performance (Liu et al., 2023).
 Figure 2.
@@
 In Part 2, GORAG performs adaptive graph retrieval and outputs the candidate labels, which is a subset of the original target label set.
-In Part 3, GORAG first classifies texts into one of the candidate labels, and then applies online indexing to update the graph with newly extracted nodes from query text.
+In Part 3, GORAG first classifies texts into one of the candidate labels, and then applies online graph indexing to update the graph with query text extracted entities that not yet existed in the graph.
 2.2.2. Long Context RAG models
-Recently, Large Language Model (LLM)-based models have undergone rapid development (Fernandez et al., 2023; Amer-Yahia et al., 2023; Zhang et al., 2024c; Miao et al., 2024; Snyder et al., 2024; Zhong et al., 2024; Yan et al., 2024; Chen et al., 2024b; Ding et al., 2024) and have been successfully adapted to various data mining tasks, including semantic parsing (Drozdov et al., 2022; Mekala et al., [n.d.]; Zhong et al., 2024; Lei et al., 2024; Sinha et al., 2024; Zhou et al., 2024), spatial-temporal mining (Zhang et al., 2024d; Li et al., 2024b; Yuan et al., 2024; Liu et al., 2024b; Wang et al., 2024a; Ning and Liu, 2024), and graph mining (Chen et al., 2024b; Tang et al., 2024; Cai et al., 2024; Jiang et al., 2024b; Cheng et al., 2024; Chen et al., 2024a; Fang et al., 2024).
-Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), making them originally suitable for dynamic text classification tasks.
-However, the lack of fine-tuning can lead LLMs to generate incorrect answers, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023).
-To mitigate hallucinations, researchers try to retrieve text documents outside of LLM as side information to help the LLM inference, namely Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Hu et al., 2024; Wu et al., 2024a; Cai et al., 2024; Oosterhuis et al., 2024; Yadav et al., 2024; Che et al., 2024).
-However, the retrieved contents from these models remain unstructured and can be lengthy, which impedes the efficiency and effectiveness of LLMs, leading to lost-in-the-middle issues (Liu et al., 2024a).
-2.2.3. Compression-based RAG models
-To address the issue suffered by Long Context RAG models, compression-based RAG models were proposed (Peng et al., 2024; Zhang et al., 2024b; Wu et al., 2024b; Li et al., 2024a; Zhang et al., 2024a; Jiang et al., 2024a; Guo et al., 2024; Gutiérrez et al., 2024), these models try to compress the side information to reduce the LLM input length.
-Based on how they compress model inputs, these models can be classified into prompt compressor models, and graph-based RAG models.
-On the one hand, prompt compressor models (Pan et al., 2024; Jiang et al., 2024a), apply LLM’s generation perplexity on to filter out un-important tokens in the model input.
-However, these models retrieve side information from each document individually, ignoring the inter-document correlations (Edge et al., 2024; Guo et al., 2024).
-On the other hand, graph-based RAG models (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) convert text chunks from the side information document as graph nodes, and the relation between these text chunks as graph edges.
-Based on the pre-defined retrieval threshold, they then retrieve a limited number of graph nodes and edges to represent the long document.
-However, the graph-based RAG models usually consider each graph edge uniformly and ignore the varying confidence and importance of relations between text chunks.
+Recently, Large Language Model (LLM)-based models have undergone rapid development (Zhang et al., 2024b; Snyder et al., 2024; Zhong et al., 2024; Yan et al., 2024; Chen et al., 2024a; Ding et al., 2024).
+Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024a), making them originally suitable for few-shot tasks.
+However, the lack of fine-tuning can lead LLMs to generate unfaithful responses, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023).
+In web for good applications such as hate speech detection, hallucinations can lead to failure to detect harmful responses, making these harmful contents being spread through internet.
+To mitigate hallucinations, researchers try to retrieve text documents outside of LLM as side information to help the LLM inference, namely Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Hu et al., 2024; Wu et al., 2024a; Cai et al., 2024; Oosterhuis et al., 2024; Yadav et al., 2024; Che et al., 2024; Pan et al., 2024; Jiang et al., 2024).
+Among them, prompt compressor models (Pan et al., 2024; Jiang et al., 2024), apply LLM’s generation perplexity on to filter out un-important tokens in the model input.
+2.2.3. Graph-based RAG models
+However, the retrieved contents of Long Context RAG models remain unstructured and lengthy, making the important fine-grained entity-level information be overwhelmed by noisy contexts (Edge et al., 2024; Guo et al., 2025), which still impedes the effectiveness or efficiency of LLM classification, leading to lost-in-the-middle issues (Liu et al., 2024b) for cheaper open-sourced LLMs, or high token cost for proprietary LLMs (Hurst et al., 2024).
+To address these issues, Graph-based RAG models were proposed (Peng et al., 2024; Zhang et al., 2024a; Wu et al., 2024b; Zhang et al., 2024c; Guo et al., 2025; Jimenez Gutierrez et al., 2024; gutiérrez2025ragmemory), these models convert text chunks from the side information document as graph nodes, and the relation between these text chunks as graph edges.
+These models then retrieve graph nodes and edges as the LLM classification context, better preserving more fine-grained, entity-level information.
+However, current Graph-based RAG models usually consider each graph edge uniformly and ignore the varying confidence and importance of relations between text chunks.
 Also, as the optimal retrieval threshold can vary across different data samples, their globally fixed threshold selected by humans may be suboptimal for the entire dataset.
-In this paper, we proposed a novel compression-based graph-RAG model, namely GORAG, which first constructs a weighted graph to model the inter-document correlations and varying confidence and importance of relations, then it achieves adaptive retrieval on this graph, avoiding any human-defined thresholds.
 3. Methodology
 3.1. Framework Overview
-To address the aforementioned uniform-indexing, threshold-dependent, and narrow-source issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph.
-As shown in Figure 4, GORAG consists of three core parts, i.e., index graph construction, graph retrieval, and text classification and online Index.
-Part 1: Graph Construction.
-It targets to construct the weighted graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ based on the labeled data $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ at the $r$-th round.
-An example the weighted graph created on WOS dataset is shown in Figure 3, graph formed like this will be used to provide retrieved-augmented information as context for query texts, enabling LLMs to better understand query texts and accurately predict their labels.
-Specifically, under each new data $\mathcal{D}_{new}^{r}$ at the $r$-th round, we first extract keywords $\mathcal{K}_{new}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$ from the labeled training texts $\mathcal{T}^{r}_{new}=\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$.
-Then, we will assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}_{new}^{r}$ and their respective text’s label $y\in\mathcal{Y}^{r}_{new}$.
-We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance with in the text corpus and relatedness to the label $y$.
-At last, we merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$.
+To address the aforementioned Uniform-indexing, Non-adaptive Retrieval, and Narrow-source issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph.
+As shown in Figure 2, GORAG consists of three core parts, i.e., Offline Graph Indexing, Graph Retrieval, and Classification & Online Graph Indexing.
+Part 1: Offline Graph Indexing.
+It targets to construct the weighted graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ based on the labeled data $\mathcal{D}^{r}=\cup_{y_{i}\in\mathcal{Y}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ at the $r$-th round.
+An example weighted graph created is shown in Figure 2, this graph will be used to provide retrieved-augmented information as contexts for query texts, enabling LLMs to better understand query texts and accurately classify them.
+Specifically, with the new labeled data $\mathcal{D}^{r}$ at the $r$-th round, we first extract keywords $\mathcal{K}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}^{r}|}$.
+Then, we will assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}^{r}$ and their respective text’s label $y\in\mathcal{Y}^{r}$, where $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}^{i}_{new}$ denotes the union label set of all new labels $\mathcal{Y}^{i}_{new}$ from round $i\in\{1,2,\dots,r\}$.
+We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance within the text corpus and keyword relatedness to the label $y$.
+At last, we merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$.
 More details can refer to Section 3.2.
 Part 2: Graph Retrieval.
-The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the query text $t$ to the constructed weighted graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all these keywords.
-From this minimum-cost spanning tree (MST), the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$.
-As this MST solely depends on the graph itself and the extracted keywords, we can get rid of any human-defined thresholds and achieve adaptive retrieval.
+The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the query text $t$ to the constructed weighted graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all mapped keywords in $\mathcal{V}^{t}$.
+From this minimum-cost spanning tree (MST), the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ at the current round $r$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$ at the current round $r$.
+As this MST solely depends on the graph itself and the text extracted keywords, we can eliminate any human-defined thresholds and achieve adaptive retrieval.
 For more details, please refer to Section 3.3.
-Part 3: Text Classification and Online Index.
-In Text Classification part, we first classify the query text with a LLM.
-This process create LLM input with the query text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the weighted graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$ if text formed label names are available.
+Part 3: Classification & Online Graph Indexing.
+In the Classification part, we create LLM input with the query text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the weighted graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$ if available.
 The LLM will carry out classification with this input.
-After classification, the online indexing procedure dynamically indexes keywords, denoted as $\mathcal{V}_{\text{notexist}}^{t}$, which are extracted from the query text $t$ but are not in the existing graph $\mathcal{G}^{r}$.
-We then integrate these keywords into the graph to further enrich its structure, with edge weights assigned based on their importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the query text $t$.
-It enhances the model’s ability to make more accurate predictions.
+After classification, the online graph indexing procedure dynamically adds keywords $\mathcal{V}_{notexist}^{t}$, which are extracted from the query text $t$ but are not in the existing graph $\mathcal{G}^{r}$, as new graph nodes to further enrich the graph.
+We then assign edge weights based on these newly added keywords’ importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the query text $t$.
+By penalizing the edges between nodes and their semantically irrelevant labels, GORAG can reduce the error introduced by inaccurate classification.
+The Online Graph Indexing procedure improves the model classification ability by updating the graph in real-time with new keywords.
 For more details, please refer to Section 3.4.
-3.2. Part 1: Graph Construction
+Due to space limitations, we put the time and space complexity analysis of GORAG to Appendix D.
+3.2. Part 1: Offline Graph Indexing
 In this subsection, we introduce the graph construction procedure of GORAG, which includes a novel edge-weighting mechanism to address the uniform indexing issue.
-The pseudo code of graph construction is shown in Algorithm 1 in Appendix Appendix C.
-For each round $r$, given its respective labeled training texts $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}_{new}$ from $\mathcal{D}_{new}^{r}$.
-Specifically, it uses the LLM model with an extraction instruction prompt $p_{extract}=$Please extract some keywords from the following passage” for this extaction.
-Then, we can get all text keywords $\mathcal{K}^{r}_{new}$ based on texts $\mathcal{D}^{r}_{new}$ as follows.
-(2) $\mathcal{K}^{r}_{new}=\bigcup_{t\in\mathcal{D}^{r}_{new}}LLM(p_{extract},t).\vspace{-5px}$
-These keywords are served as graph nodes, we denote them as keyword nodes.
-Also, GORAG incorporats new candidate labels $\mathcal{Y}^{r}_{new}$ at the $r$-th round as graph nodes, denoted as label nodes.
+The pseudo code of graph construction is shown in Algorithm 1 in Appendix C.
+For each round $r$, given its respective labeled training texts $\mathcal{D}^{r}=\cup_{y_{i}\in\mathcal{Y}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}$ from $\mathcal{D}^{r}$.
+Specifically, it calculates the normalized TFIDF score (69) as the correlation score $CS(v,t)$ for texts’ each word and phrase with less than 3 words (Wang et al., 2025).
+(2) $\vskip-4.015ptCS(v,t)=\frac{count(v,t)}{|t|log(|\mathcal{T}^{r}|+1)}\times log\frac{|\mathcal{T}^{r}|+1}{|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|+1},$
+where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}^{i}$ denotes all training and query texts seen so far, $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$.
+Here, we will extract words or phrases with $CS(v,t)>\tau$ as our keyword.
+We set $\tau=0.5$ following the keyword extraction setting in (Wang et al., 2025).
+These keywords serve as graph nodes, and we denote them as keyword nodes.
+Also, GORAG incorporates new labels $\mathcal{Y}^{r}_{new}=\mathcal{Y}^{r}\setminus\mathcal{Y}^{r-1}$ at the $r$-th round as graph nodes, denoted as label nodes, where $\mathcal{Y}^{r}$ is the total label set at round $r$.
 Hence, the graph node set $\mathcal{V}^{r}_{new}$ at the $r$-th round can be obtained as follows.
-(3) $\mathcal{V}_{new}^{r}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}.$
-Then, GORAG add edges between each keyword node $v\in\mathcal{K}_{{new}}^{r}$ to its corresponding label node $y\in\mathcal{Y}_{{new}}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
-Next, considering the keywords $\mathcal{K}_{new}^{r}$ from the labeled text are not uniformly related to the text’s label $\mathcal{Y}_{new}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label edge.
-This weight can be regarded as the correlation between keywords and each label.
-Firstly, we apply the normalized TF-IDF score (Sammut and Webb, 2010) as our correlation score $CS(v,t)$ to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}_{new}^{r}$ w.r.t.
-the respective text $t$ it extracted from:
-(4) $\vspace{-5px}CS(v,t)=\frac{count(v,t)}{|t|}\times log\frac{|\mathcal{T}^{r}|}{1+|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|},$
-where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}_{new}^{i}$ denotes all training and query texts seen so far, $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$.
+(3) $\vskip-5.01874pt\mathcal{V}_{new}^{r}=\mathcal{K}^{r}\cup\mathcal{Y}^{r}_{new}.$
+Then, GORAG adds edges between each keyword node $v\in\mathcal{K}^{r}$ to its corresponding label node $y\in\mathcal{Y}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
+Next, considering the keywords $\mathcal{K}^{r}$ from the labeled text are not uniformly related to the text’s label in $\mathcal{Y}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label edge between any two nodes $v$ and $y$.
+This weight can reflect the correlation between keywords and each label.
+Firstly, we apply the correlation score $CS(v,t)$ assigned by Equation (2) to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}^{r}$ w.r.t.
+the text $t$ it extracted from.
 As the keyword $v$ can be extracted from multiple texts with different labels in different rounds, the final edge weight $w_{v,y}^{r}$ of edge $e_{v,y}$ at the $r$-th round is calculated as an average of all weights from all seen texts with label $y$:
-(5) $\vspace{-3px}w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,y,v}}1-CS(v,t_{j})}{|\mathcal{T}^{r,y,v}|},$
-where $\mathcal{T}^{r,y,v}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text that contains keyword $v$ and labeled $y$.
-We denote the generated weighted graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where $\mathcal{V}^{r}_{new}$, $\mathcal{E}^{r}_{new}$, and $\mathcal{W}^{r}_{new}$ denotes the node, edge and edge weight set respectively, an example of the generated graph is shown in Figure 3.
+(4) $w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,v,y}}1-CS(v,t_{j})}{|\mathcal{T}^{r,v,y}|},$
+where $\mathcal{T}^{r,v,y}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text that contains keyword $v$ and labeled $y$.
+We denote the generated weighted graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where $\mathcal{V}^{r}_{new}$, $\mathcal{E}^{r}_{new}$, and $\mathcal{W}^{r}_{new}$ denotes the node, edge and edge weight set respectively.
 Lastly, $\mathcal{G}^{r}_{new}$ will be merged into the graph from previous rounds $\mathcal{G}^{r-1}$ to form the full graph $\mathcal{G}^{r}$ at round $r$:
@@
 Particularly, we define $\mathcal{G}^{0}=\emptyset$ and $r\geq 1$.
-Also, to guarantee the graph connectivity of the resulting graph, we add edges between every new label node $y_{n}\in\mathcal{Y}_{new}^{r}$ and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{y_{n},y_{o}}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is defined as the weighted average of all edge weights that link keyword nodes with label node $y_{n}$ or $y_{o}$:
-(7) $\vspace{-5px}\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}\vspace{-5px}$
-(8) $\vspace{-3px}w_{y_{n},y_{o}}^{r}=\frac{1}{2}\big{(}\frac{\sum_{v\in\mathcal{M}_{y_{n}}^{r}}w_{v,y_{n}}^{r}}{2\times|\mathcal{M}_{y_{n}}^{r}|}+\frac{\sum_{v\in\mathcal{M}_{y_{o}}^{r}}w_{v,y_{o}}^{r}}{2\times|\mathcal{M}_{y_{o}}^{r}|}\big{)}$
+Also, to guarantee the graph connectivity of the resulting graph, we add edges between every new label node $y_{n}\in\mathcal{Y}_{new}^{r}$, and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{y_{n},y_{o}}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is defined as the weighted average of all edge weights that link keyword nodes with label node $y_{n}$ or $y_{o}$:
+(6) $\vskip-5.01874pt\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}\vskip-5.01874pt$
+(7) $\vskip-3.01125ptw_{y_{n},y_{o}}^{r}=\frac{1}{2}\Big(\frac{\sum_{v\in\mathcal{M}_{y_{n}}^{r}}w_{v,y_{n}}^{r}}{2\times|\mathcal{M}_{y_{n}}^{r}|}+\frac{\sum_{v\in\mathcal{M}_{y_{o}}^{r}}w_{v,y_{o}}^{r}}{2\times|\mathcal{M}_{y_{o}}^{r}|}\Big),$
 where $\mathcal{N}^{r}(y)$ denote the neighbor nodes of label node $y$ in $\mathcal{G}^{r}$.
@@
 With the graph constructed in Part 1, GORAG can adaptively retrieve a set of candidate class labels with keywords extracted from query texts without any human-defined thresholds, addressing the threshold dependent issue.
-To begin with, GORAG extract keywords $\mathcal{V}_{test}^{t}$ for each query text $t\in\mathcal{T}_{test}$ in the same manner with Equation (2), then, $\mathcal{V}_{test}^{t}$ would be splitted into two subsets:
+To begin with, GORAG extracts keywords $\mathcal{V}_{test}^{t}$ for each query text $t\in\mathcal{T}_{test}$ in the same manner with Equation (2), then, $\mathcal{V}_{test}^{t}$ would be split into two subsets:
 (8) $\mathcal{V}_{exist}^{t}\cup\mathcal{V}_{notexist}^{t}=\mathcal{V}_{test}^{t},\ \mathcal{V}_{exist}^{t}\cap\mathcal{V}_{notexist}^{t}=\emptyset,$
@@
 Later, $\mathcal{V}_{exist}^{t}$ would be applied for achieving adaptive retrieval, and $\mathcal{V}_{notexist}^{t}$ would be applied for online indexing to further enrich the constructed graph (Further illustrated in Section 3.4).
-To achieve the adaptive retrieval, GORAG trys to find the minimum cost spanning tree (MST) that contain all keyword nodes within $\mathcal{V}_{exist}^{t}$, and then retrieve all label nodes within the generated MST as candidate labels.
-The intuition behind this approach is that a MST spans the entire graph to cover all given nodes with the smallest possible spanning cost.
+To achieve the adaptive retrieval, GORAG tries to find the minimum cost spanning tree (MST) that contain all keyword nodes within $\mathcal{V}_{exist}^{t}$, and then retrieves all label nodes within the generated MST as candidate labels.
+The intuition behind this approach is that an MST spans the entire graph to cover all given nodes with the smallest possible spanning cost.
 Consequently, label nodes within the generated MST can be considered important for demonstrating the features of the given keyword node set.
@@
 Table 1.
-Statistics of the WOS, Reuters, and IFS-Rel datasets, where we divide their original labels into multiple rounds.
-We achieve a balanced testing set on the Reuters and IFS-Rel datasets by assigning each label 10 and 40 testing samples, respectively.
+Statistics of the tested datasets, where we divide their original labels into multiple rounds.
+We achieve a balanced testing set on the IFS-Rel datasets by assigning each label 40 testing samples.
 Dataset Avg. Text Token R1 R2 R3 R4 Total
 Testing data Label # Testing data Label # Testing data Label # Testing data Label # Testing data
+CAD (Vidgen et al., 2021) 154 240 4 60 3 286 3 231 3 817
 WOS (Kowsari et al., 2017) 200 2,417 32 2,842 53 2,251 30 1,886 18 9,396
-Reuters (Bao et al., 2020) 168 80 8 80 8 80 8 70 7 310
 IFS-Rel (Xia et al., 2021) 105 640 16 640 16 640 16 640 16 2,560
@@
 $\min_{e^{r}_{u.v}\in\mathcal{E}^{r}_{t}}w^{r}_{u,v}$
-(10) $s.t.$ $v\in\mathcal{V}^{r}_{t},\forall v\in\mathcal{V}_{exist}^{t}$
-Then, since the subgraph nodes $\mathcal{V}^{r}_{t}$ contains both keyword nodes and label nodes, we take the label nodes ${\mathcal{Y}}^{r}_{t}\in\mathcal{V}^{r}_{t}$ in the gernerated sub-graph as our target candidate nodes for the text $t$.
-Since this problem is the NP-hard (Su et al., 2020), it is infeasible to obtain the optimal result in polynomial time.
-Therefore, to solve this problem, we propose a greedy algorithm based on the Melhorn’s algorithm (Mehlhorn, 1988a).
-Our algorithm generates the candidate label set ${\mathcal{Y}}^{r}_{t}\subseteq\mathcal{Y}^{r}$ for text $t$ at the $r$-th round.
-The pseudo code of our algorithm is shown in Appendix C, following the Melhorn’s algorithm, our algorithm can also achieve 2-approximate.
-Firstly, we calculate the shortest path between each keyword node $v$ in $\mathcal{V}_{exist}^{t}$ to all other nodes in $\mathcal{G}^{r}$ by calculating the minimum spanning tree $MST$ of $\mathcal{G}^{r}$ Here, the keyword nodes $v$ in $\mathcal{V}_{exist}^{t}$ are served as terminal nodes that determines the final generated candidate labels w.r.t.
-to the weighted graph $\mathcal{G}^{r}$ Secondly, we create a new auxiliary graph $H$ where the edges represent the shortest paths between the closest terminal nodes.
+(9) $s.t.\$$v\in\mathcal{V}^{r}_{t},\forall v\in\mathcal{V}_{exist}^{t}$
+Then, since the subgraph node $\mathcal{V}^{r}_{t}$ set contains both keyword nodes and label nodes, we define the candidate label set $\hat{\mathcal{Y}}^{r}_{t}\in\mathcal{V}^{r}_{t}$ as the label nodes in the generated sub-graph for the text $t$ at round $r$.
+Due to space limitations, we put the proof of the Adaptive Candidate Label Generation Problem to be NP-hard to Appendix B.
+As the problem is NP-hard, it is infeasible to obtain the optimal result in polynomial time.
+Therefore, to solve this problem, we propose a greedy algorithm based on the Melhorn’s algorithm (Mehlhorn, 1988).
+Our algorithm generates the candidate label set $\hat{\mathcal{Y}}^{r}_{t}\subseteq\mathcal{Y}^{r}$ for text $t$ at the $r$-th round.
+The pseudo code of our algorithm is shown in Algorithm 2 in Appendix C , following the Melhorn’s algorithm, our algorithm can also achieve 2-Approximate.
+Firstly, we calculate the shortest path between each keyword node $v$ in $\mathcal{V}_{exist}^{t}$ to all other nodes in $\mathcal{G}^{r}$ by calculating the minimum spanning tree $MST$ of $\mathcal{G}^{r}$ Here, the keyword nodes $v$ in $\mathcal{V}_{exist}^{t}$ serve as terminal nodes that determines the final generated candidate labels w.r.t.
+to the weighted graph $\mathcal{G}^{r}$.
+Secondly, we create a new auxiliary graph $H$ where the edges represent the shortest paths between the closest terminal nodes.
 Thirdly, we construct the minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$ , and then add the shortest paths between each two nodes in $MST^{\prime}$ to the Steiner Tree $ST$.
-Lastly our candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ for text $t$ are calculated as the interselect of $ST$ and all target labels $\mathcal{Y}^{r}$ at round $r$: $ST\cap\mathcal{Y}^{r}$.
-Please refer to Appendix C for more details.
-3.4. Part 3: Text Classification and Online Index
-In this subsection, we introduce GORAG’s classification part, where it futher enrich the weighted graph by adding query text information with the online indexing mechanism, addressing the narrow source issue.
+Lastly our candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ for text $t$ are calculated as the intersection of $ST$ and all target labels $\mathcal{Y}^{r}$ at round $r$: $ST\cap\mathcal{Y}^{r}$.
+3.4. Part 3: Classification & Online Indexing
+In this subsection, we first introduce GORAG’s classification and online graph indexing procedure, where it classifies the texts and further enriches the weighted graph by adding query text information to the created graph, addressing the narrow source issue.
 To begin with, we introduce how GORAG performs text classification based the LLM.
-Specifically, given each unlabled query text $t$, GORAG utilizes LLM to predict its class label $y^{*}$ by constructing an input prompt $c_{t}$ as follows:
+Specifically, given each unlabeled query text $t$, GORAG utilizes LLM to predict its class label $y^{*}$ by constructing an input prompt $c_{t}$ as follows:
 (10) $c_{t}=Concat(t,\mathcal{K}^{t},\hat{\mathcal{Y}}^{r}_{t},\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}).$
-$c_{t}$ can be considered as the text concatenation of the extracted keywords $\mathcal{K}^{t}$ from the text $t$, and the candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ obtained by Algorithm 2.
+$c_{t}$ can be considered as the text concatenation of the extracted keywords $\mathcal{K}^{t}$ from the text $t$, and the candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ obtained by Algorithm 2 in Appendix C, we omit the more complex input structures as they rarely improve performance, but lead to higher costs (Zhou et al., 2023; Liu et al., 2024b).
 Also, the $\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}=\{\mathcal{K}_{y_{i}}\}_{i=1}^{|\hat{\mathcal{Y}}^{r}_{t}|}$ and $\mathcal{K}_{y_{i}}$ are the representative keywords of each label $y_{i}\in\hat{\mathcal{Y}}^{r}_{t}$.
 (11) $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r}),$
-which were generated by LLM with the label’s semantic label name if avaiable. Next, with a classification instruction prompt $p_{classify}$.
-(13) $\vspace{-5px}y^{*}_{t}=LLM(p_{classify},c_{t}),$
-the LLM would try to select the best-suited label $y^{*}_{t}\in\mathcal{Y}^{t}$ to annotate the text $t$.
+which were generated by LLM with the label’s semantic label name if available. Next, with a classification instruction prompt $p_{classify}$.
+(12) $y^{*}_{t}=LLM(p_{classify},c_{t}),$
+the LLM would try to select the best-suited label $y^{*}_{t}\in\mathcal{Y}^{r}$ to annotate the text $t$.
 To fully leverage the text-extracted keywords, GORAG utilizes an online indexing mechanism to incrementally update keywords that do not yet exist in the weighted graph $\mathcal{G}^{r}$ at the $r$-th round to $\mathcal{G}^{r}$ based on the text $t$’s predicted label $y^{*}_{t}$.
 To be specific, each keyword node $v\in\mathcal{V}_{notexist}^{t}$ would be added to the original graph’s node set $\mathcal{V}^{r}$ and be assigned with an edge $e_{v,y}$ connecting it with the text $t$’s predicted label $y_{t}^{*}$:
-(14) $\vspace{-5px}\mathcal{V}^{r}=\mathcal{V}^{r}\cup\mathcal{V}_{notexist}^{t},\ \mathcal{E}^{r}=\mathcal{E}^{r}\cup\mathcal{E}_{oi}^{t},$
+(13) $\vskip-3.01125pt\mathcal{V}^{r}=\mathcal{V}^{r}\cup\mathcal{V}_{notexist}^{t},\ \mathcal{E}^{r}=\mathcal{E}^{r}\cup\mathcal{E}_{oi}^{t},$
 where $\mathcal{E}_{oi}^{t}=\{e_{v,y_{t}^{*}}|v\in\mathcal{V}_{notexist}^{t}\}$ denotes the set of all newly assigned edges $e_{v,y}$ between keyword node $v$ and its predicted label $y_{t}^{*}$, for these newly assigned edges, their weight is calculated with the edge weighting mechanism illustrated in Equation (4).
-This online indexing mechanism has two purposes.
-Firstly, it incrementally enriches the weighted graph $\mathcal{G}^{r}$ by incorporating new, previously unseen keywords, thus expanding the graph’s vocabulary and improving its adaptability.
-Secondly, by linking these new keywords to their predicted labels with weighted edges, this mechanism captures their relevance and context more effectively, enabling better utilization of the graph for future predictions.
+The edge weights can reflect the semantic relevance of keywords to labels, and reduce the noise introduced by inaccurate LLM classification.
 4. Experiments
-In this section, we present the experimental evaluation of our framework GORAG on the DFSTC task.
-We compare the performance of GORAG against six effective baselines spanning three technical categories, using three datasets with distinct characteristics.
-Specifically, we first outline the experimental setup in Section 4.1, including details on datasets, and evaluation metrics.
-Due to the space limitation, we put the details of baselines, and hyperparameter configurations in Appendix Section E.
-Next, we report the experimental results, focusing on both effectiveness and efficiency evaluations, in Section 4.2.
-Finally, we conduct an ablation study and a case study, presented in Section 4.5 and Appendix Section E.5.
+Table 2. Classification accuracy on CAD, WOS, and IFS-Rel benchmarks.
+Dataset Category Model Round 1 Round 2 Round 3 Round 4
+1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
+CAD NN-based Entailment (Xia et al., 2021) 0.2125 0.2000 0.1208 0.1700 0.1060 0.1634 0.1177 0.0836 0.0785 0.0857 0.0700 0.0529
+Long Context RAG NaiveRAG (Gao et al., 2023) 0.1875 0.1375 0.0958 0.1500 0.1100 0.0766 0.0802 0.0563 0.0392 0.0600 0.0405 0.0281
+LongLLMLingua (Jiang et al., 2024) 0.1667 0.2417 0.2417 0.1709 0.2167 0.2200 0.1370 0.1587 0.1552 0.1092 0.1359 0.1285
+Graph-based RAG GraphRAG (Edge et al., 2024) 0.3958 0.3625 0.3292 0.3167 0.2933 0.2833 0.1587 0.1621 0.1706 0.1200 0.1236 0.1322
+LightRAG (Guo et al., 2025) 0.3875 0.3583 0.3373 0.3133 0.2867 0.2867 0.1638 0.1502 0.1571 0.1175 0.1102 0.1188
+HippoRAG2 (gutiérrez2025ragmemory) 0.3667 0.2250 0.1792 0.2953 0.2000 0.1458 0.1820 0.1032 0.0746 0.1435 0.0668 0.0535
+GORAG 0.6125 0.6125 0.6125 0.5867 0.5833 0.5882 0.3993 0.4010 0.4001 0.3133 0.3146 03139
+WOS NN-based Entailment (Xia et al., 2021) 0.3695 0.3823 0.4187 0.3994 0.4471 0.4222 0.4510 0.4857 0.4787 0.4030 0.4387 0.4442
+Long Context RAG NaiveRAG (Gao et al., 2023) 0.3885 0.3904 0.3897 0.2267 0.2154 0.2187 0.1821 0.1475 0.1799 0.1653 0.1556 0.1649
+LongLLMLingua (Jiang et al., 2024) 0.3806 0.3823 0.3901 0.2155 0.2202 0.2198 0.1770 0.1567 0.1608 0.1468 0.1382 0.1493
+Graph-based RAG GraphRAG (Edge et al., 2024) 0.3852 0.3897 0.3906 0.2213 0.2197 0.2219 0.1816 0.1770 0.1786 0.1641 0.1634 0.1625
+LightRAG (Guo et al., 2025) 0.3930 0.3806 0.3815 0.2202 0.2216 0.2145 0.1743 0.1767 0.1799 0.1625 0.1626 0.1632
+HippoRAG2 (gutiérrez2025ragmemory) 0.4133 0.4245 0.4208 0.2552 0.2656 0.2666 0.2205 0.2210 0.2242 0.1906 0.1860 0.1915
+GORAG 0.4866 0.4990 0.5134 0.4790 0.5109 0.5284 0.4736 0.4996 0.5234 0.4420 0.4717 0.4929
+IFS-Rel NN-based Entailment (Xia et al., 2021) 0.3391 0.6046 0.5859 0.4008 0.5516 0.5828 0.3061 0.4703 0.5193 0.2971 0.4395 0.4895
+Long Context RAG NaiveRAG (Gao et al., 2023) 0.7343 0.7344 0.7406 0.5394 0.5406 0.5469 0.4719 0.4719 0.4688 0.5233 0.4938 0.5000
+LongLLMLingua (Jiang et al., 2024) 0.8750 0.8906 0.9219 0.5734 0.5922 0.5806 0.4703 0.4741 0.5225 0.5819 0.4903 0.5038
+Graph-based RAG GraphRAG (Edge et al., 2024) 0.6406 0.7813 0.8125 0.4406 0.4688 0.4719 0.3828 0.3734 0.3141 0.4297 0.4422 0.3922
+LightRAG (Guo et al., 2025) 0.7578 0.8281 0.8594 0.4906 0.5155 0.5156 0.3802 0.4010 0.4427 0.4531 0.4804 0.4610
+HippoRAG2 (gutiérrez2025ragmemory) 0.7313 0.7375 0.7406 0.4875 0.5016 0.4938 0.4297 0.4328 0.4438 0.4984 0.5008 0.4953
+GORAG 0.9688 0.9688 0.9688 0.5938 0.6172 0.6250 0.4844 0.4948 0.5344 0.6165 0.4988 0.5469
+In this section, we evaluate GORAG’s performance on the DFSTC task.
+Section 4.1 introduces the experimental setup, while Section 4.2 presents results on effectiveness and efficiency.
+Section 4.3 covers ablation studies, and Section 4.4 discusses generalization.
+Due to space limitation, we put experiment details, such as hardware settings and case study in Appendix E.
 4.1. Experiment Settings
@@
 4.1.2. Datasets
-We select the Web-Of-Science (WOS) dataset (Kowsari et al., 2017) and the Reuters few-shot text classification dataset by (Bao et al., 2020) for evaluating GORAG’s performance.
-We chose these datasets because they have a considerable amount of classes and text tokens, which is complex and can reflect the real-world scenario.
-As shown in Table 1, texts from these datasets can contain multiple sentences and have average tokens ranging from 105 to 200 per text.
-The IFS-Rel dataset was extended from its original version in (Xia et al., 2021), which only has 32 labels in total and an average of 25 tokens per text.
-Table 2. Classification accuracy on WOS, Reuters, and IFS-Rel dataset.
-Dataset Category Model Round
-R1 R2 R3 R4
-1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
-WOS NN-based Entailment (Xia et al., 2021) 0.3695 0.3823 0.4187 0.3994 0.4471 0.4222 0.4510 0.4857 0.4787 0.4030 0.4387 0.4442
-Long Context RAG NaiveRAG (Gao et al., 2023) 0.3885 0.3904 0.3897 0.2267 0.2154 0.2187 0.1821 0.1475 0.1799 0.1653 0.1556 0.1649
-Propositionizer (Chen et al., 2023) 0.1241 0.1306 0.1297 0.1074 0.1421 0.1303 0.1771 0.1645 0.1603 0.1611 0.1637 0.1656
-Compression RAG LongLLMLingua (Jiang et al., 2024a) 0.3806 0.3823 0.3901 0.2155 0.2202 0.2198 0.1770 0.1567 0.1608 0.1468 0.1382 0.1493
-GraphRAG (Edge et al., 2024) 0.3852 0.3897 0.3906 0.2213 0.2197 0.2219 0.1816 0.1770 0.1786 0.1641 0.1634 0.1625
-LightRAG (Guo et al., 2024) 0.3930 0.3806 0.3815 0.2202 0.2216 0.2145 0.1743 0.1767 0.1799 0.1625 0.1626 0.1632
-GORAG 0.4866 0.4990 0.5134 0.4790 0.5109 0.5284 0.4736 0.4996 0.5234 0.4380 0.4717 0.4929
-Reuters NN-based Entailment (Xia et al., 2021) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
-Long Context RAG NaiveRAG (Gao et al., 2023) 0.0000 0.0000 0.0375 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
-Propositionizer (Chen et al., 2023) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
-Compression-RAG LongLLMLingua (Jiang et al., 2024a) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0258 0.0065 0.0085
-GraphRAG (Edge et al., 2024) 0.1375 0.1625 0.1000 0.0688 0.0813 0.0500 0.0375 0.0417 0.0417 0.0291 0.0375 0.0375
-LightRAG (Guo et al., 2024) 0.0500 0.1375 0.1125 0.0250 0.0813 0.0563 0.0333 0.0333 0.0333 0.0125 0.0333 0.0208
-GORAG 0.0875 0.0875 0.1000 0.1750 0.1688 0.1438 0.1667 0.1958 0.2167 0.1667 0.2516 0.2193
-IFS-Rel NN-based Entailment (Xia et al., 2021) 0.3391 0.6046 0.5859 0.4008 0.5516 0.5828 0.3061 0.4703 0.5193 0.2971 0.4395 0.4895
-Long Context RAG NaiveRAG (Gao et al., 2023) 0.4844 0.4719 0.4828 0.3711 0.3555 0.3656 0.2708 0.2750 0.2625 0.2516 0.2465 0.2527
-Propositionizer (Chen et al., 2023) 0.1719 0.1828 0.1938 0.1602 0.1727 0.1750 0.1792 0.1552 0.1770 0.1574 0.1688 0.1664
-Compression-RAG LongLLMLingua (Jiang et al., 2024a) 0.4703 0.5203 0.4297 0.3458 0.3055 0.3733 0.2452 0.2206 0.2713 0.2154 0.2029 0.1997
-GraphRAG (Edge et al., 2024) 0.4797 0.4656 0.4734 0.3672 0.3641 0.3641 0.2740 0.2776 0.2708 0.2453 0.2465 0.2426
-LightRAG (Guo et al., 2024) 0.4484 0.4234 0.4656 0.3531 0.3477 0.3523 0.2760 0.2760 0.2682 0.2340 0.2367 0.2391
-GORAG 0.4875 0.5266 0.5313 0.4195 0.3781 0.4374 0.2953 0.2974 0.3328 0.2973 0.2571 0.3034
-Table 3. The size of the constructed weighted graph in each round.
-Dataset Model Round
-R1 R2 R3 R4 After R4
-Node # Edge # Node # Edge # Node # Edge # Node # Edge # Node # Edge #
-WOS GORAG offline 1,149 1,405 3,681 4,587 4,947 6,204 5,438 6,879 5,438 6,879
-GORAG 1,149 1,405 12,108 12,245 25,357 26,290 34,521 35,919 44,283 45,973
-Reuters GORAG offline 117 119 192 200 263 284 331 366 331 366
-GORAG 117 119 584 594 1,013 1,046 1,342 1,405 1,533 1,600
-IFS-Rel GORAG offline 189 213 242 275 372 469 454 594 454 594
-GORAG 189 213 5,299 5,503 9,417 9,934 13,293 14,074 16,705 17,486
-We split the WOS dataset into 4, 6 and 8 rounds, the Reuters and IFS-Rel datasets into 4 rounds, in order to maintain a comparable number of new label and testing data in each round.
-Due to the space limitation, we mainly display our experinents on the 4 rounds spit version of both datasets in Table 2, and the experiment result of 6 and 8 rounds are shown in Figure 5(a) and 5(b).
-Different from the WOS and IFS-Rel datasets, the labels of Reuters are in numeric formats, and their text-formed label names are not available.
-According to Table 2, this feature makes many baselines fail to carry out classification task on the Reuters dataset.
-Further statistics of these two dataset being splitted into 4 rounds are shown in Table 1.
+We select the CAD (Vidgen et al., 2021), WOS (Kowsari et al., 2017), and IFS-Rel (Xia et al., 2021) benchmark to test GORAG’s performance.
+The texts in CAD are social media posts that potentially have different kinds of hate speech and abuse.
+The texts in the IFS-Rel benchmark are segments from web news, and their classifications help detect misinformation.
+The texts from the WOS benchmark are abstracts or introductions from academic papers, as the WOS dataset contains a vast number of candidate labels and has a large word count for each text; these characteristics make it a challenging benchmark for text classification models.
+The statistics of these benchmarks are shown in Table 1.
+Due to the space limitations, we mainly display our experinents on the 4 rounds split version of all datasets in Table 2, and further experiment results of 6 and 8 rounds are shown in Figure 4.
+Ethical Disclaimer.
+This research involves sensitive content.
+All datasets are publicly available, and our study adheres to institutional ethical guidelines to ensure that no personally identifiable information is disclosed.
+Examples shown in this paper are for illustrative purposes only and do not reflect the authors’ views.
 4.1.3. Evaluation Metrics
-In this paper, aligning with previous DFSTC models (Xia et al., 2021; Meng et al., 2020), we use classification accuracy as the main evaluation metric.
-Given that LLMs can generate arbitrary outputs that may not precisely match the provided labels, we consider a classification correct only if the LLM’s output exactly matches the ground-truth label name or label number, if the model cannot give out a class within the provided target label set, we denote this classification as a hallucination.
-We further provide the macro recall score and the hallucination rate on the most complicated and imbalanced test dataset WOS in Table 8 in Appendix.
+In this paper, aligning with previous DFSTC models (Xia et al., 2021; Meng et al., 2020), we use classification accuracy as the evaluation metric.
+Given that LLMs can generate arbitrary outputs that may not precisely match the provided labels, we consider a classification correct only if the LLM’s output exactly matches the ground-truth label name or label number.
 4.1.4. Baselines
-As shown in Table 2, we compare GORAG’s performance with 7 baselines from 3 technical categories for few-shot experiments, they are NN-based Entailment (Xia et al., 2021), Long Context-based NaiveRAG (Gao et al., 2023), Propositionizer (Chen et al., 2023), and Compression-based LongLLMLingua (Jiang et al., 2024a), GraphRAG (Edge et al., 2024), and LightRAG (Guo et al., 2024).
-Further illustrations of these chosen baselines are listed in subsection E.1 in Appendix.
-For GraphRAG and NaiveRAG, we use the implementation from (nan, [n.d.]) which optimizes their original code and achieves better time efficiency while not affect the performance; For all other baselines, we use their official implementations.
-Also, long text inputs affect LLM performance and increase costs (Chen et al., 2023; Jiang et al., 2024a; Liu et al., 2024a).
-WOS and Reuters datasets, with lengthy texts and many labels, caused GPU out-of-memory (e.g., llama3-8B, 1-shot, A100-40GB).
-Thus, we compare open-source LLMs with GORAG in the 0-shot setting (Table 4).
+We compare six baselines from three types as follows. The details of selected baselines are in Appendix E.1.
+  • One NN-based model: Entailment (Xia et al., 2021), which converts the task to binary classification, and creates entailment pairs with the labeled data for training.
+  • Two Long Context RAG models: NaiveRAG (Gao et al., 2023) splits few-shot labeled data into chunks and retrieves the most query similar chunks to augment LLM classification.
+LongLLMLingua (Jiang et al., 2024) further compresses the split chunks of NaiveRAG by filtering out chunk tokens regarding their LLM generation perplexity.
+  • Three Graph-based RAG models: GraphRAG (Edge et al., 2024) extracts entities and relations from few-shot training data, and uses LLM to summarize graph neighborhoods as communities for retrieval.
+LightRAG (Guo et al., 2025) drops the community-related procedure and solely retrieves the 1-hop neighborhood of the unlabeled text mapped entities on the graph.
+HippoRAG2 (gutiérrez2025ragmemory) retrieves the Top-k candidate labels and their training texts based on their Personalized PageRank (PPR) score on the graph.
+Furthermore, as solely put the labeled text within the LLM context can greatly increase costs and affect LLM performance (Chen et al., 2023; Jiang et al., 2024; Liu et al., 2024b), we compare the performance of solely using LLM for classification with GORAG under 0-shot setting in Table 6.
+4.1.5. Hyperparameter Settings
+Following the exact setting in Entailment’s original paper (Xia et al., 2021), we train the Entailment model with the RoBERTa-large PLM for 5 epochs with a batch size of 16, a learning rate of $1\times 10^{-6}$, and the Adam optimizer (Kingma, 2014).
+For LongLLMLingua, we test the compression rate within 0.75,0.8,0.85{0.75,0.8,0.85} and select 0.8, as it achieves the best overall classification accuracy.
+For all RAG-based baselines, we set all their hyperparameter the same as in their original paper, if not further illustrated.
+For GraphRAG and LightRAG, we use their local search mode, as it achieves the highest classification accuracy on the most complex WOS dataset.
+For GraphRAG and NaiveRAG, we use the implementation from (20), which optimizes their original code and achieves better time efficiency while not affect the performance.
+For all other baselines, we use their official implementations.
 4.2. Few-shot Experiments
+(a) Time cost on WOS dataset.
+(b) Time cost on CAD dataset.
+Figure 3. Total time cost on 2 datasets under 1-shot setting.
 4.2.1. Effectiveness Evaluation
-As shown in Table 2, GORAG achieves the best classification accuracy over all four rounds with 1-shot and 5-shot labeled indexing data on WOS dataset, it surpasses all RAG-based baselines as well as the state-of-the-art model Entailment.
-Compared with Entailment, when apply 1-shot setting, GORAG achieves at most 31.6% accuracy gain at the first round.
-On Reuters dataset, GORAG achieves the best classification accuracy over the last 3 rounds with 1-shot indexing data.
-Furthermore, based on the experiment results in Table 2, we can have following observations:
-  • Firstly, NN-based Entailment fails on Reuters due to its need for text-formed label names.
-As noted in prior studies (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023), it overfits easily, with recall and hallucination issues shown in Table 8.
-Further analysis is in Appendix subsection E.3.
-  • Secondly NaiveRAG performs better than Entailment in round 1 of the WOS and IFS-Rel datasets by avoiding overfitting, but its long LLM inputs hinder its ability to adapt to dynamic label updates, leading to a significant performance drop from rounds 2 to 4.
-While Propositionizer retrieves more fine-grained results, its lack of task awareness results in consistently poor performance on DFSTC tasks.
-On the Reuters dataset, both Long Context RAG baselines fail when label names are unavailable, as the absence of semantic-rich labels prevents effective classification, especially with long LLM inputs.
-  • Thirdly, compression-based RAG models shorten LLM inputs and emphasize labels and key terms but only slightly improve performance across datasets.
+In Table 2, we apply Qwen2.5-7B backbone for CAD and WOS benchmark.
+Due to cost and API risk control (failing to respond when detecting dirty words) reasons, we only use the GPT-4o backbone on the IFS-Rel news benchmark.
+As shown in Table 2, GORAG achieves the best classification accuracy over all dynamic rounds on all benchmarks.
+Based on the experiment results, we have the following observations:
+  • Firstly, with the training data increase (R1 to R4), NN-based Entailment suffers from performance decrease on CAD datasets.
+This is because the CAD datasets have an unbalanced label distribution, as noted in prior studies (Liu et al., 2023, 2024a; Lei et al., 2023), NN-based models that apply data augmentation techniques are prone to overfit when the ground truth label distribution are unbalanced.
+  • Secondly, NaiveRAG performs better than Entailment in round 1 of the WOS and IFS-Rel datasets by avoiding overfitting, but its long inputs hinder its ability to adapt to dynamic label updates, leading to a significant performance drop from R1 to R4.
+  • Thirdly, Graph-based RAG baselines only slightly improve performance across datasets.
 They struggle with dynamic label updates due to issues like uniform indexing, threshold dependency, and narrow sources, leading to inaccurate retrievals.
-Additionally, failing to compress label sets results in lengthy LLM inputs as new labels are introduced in each round.
-  • Lastly, GORAG addresses the issues of Compression-based RAG models.
-From R1 to R4, GORAG adapts better to dynamic label updates.
-On the Reuters dataset, when label names are not available, GORAG achieves comparable or better performance than other baselines.
-According to Table 3, a vast amount of knowledge is added to the weighted graph by the online indexing mechanism, and contributes to our effectiveness.
+  • Lastly, from R1 to R4, GORAG adapts better to dynamic label updates, addresses the issues of Graph-based RAG models, and achieve better performance than all tested baselines.
+Table 3. Graph incremental trends and memory usage on IFS-Rel dataset round 1 with GPT-4o LLM.
+Object 25% 50% 75% 100%
+Nodes 1644 3238 (+1394) 4482 (+1244) 5299 (+817)
+Edges 1771 3265 (+1494) 4597 (+1332) 5503 (+906)
+Memory 13.3 KB 27.6 KB 37.0 KB 46.4 KB
 4.2.2. Efficiency Evaluation
-(a) 1-shot setting.
-(b) 5-shot setting.
-(c) Time cost on WOS dataset.
-(d) Time cost on Reuters dataset.
-Figure 5. Experiment result of WOS dataset under 1-shot and 5-shot settings for at most 8 rounds, and model Training/Indexing time cost on two datasets under 1-shot setting.
-To evaluate the efficiency of GORAG, we compare GORAG’s indexing time cost with other baselines’ training or indexing time cost on the 4 round split version of WOS and Reuters datasets, the results are shown in Figure 5(c) and 5(d).
-For both datasets, Propositionizer achieves the worst indexing efficiency.
-This is because it employs an additional fine-tuned LLM to convert each side information into propositions.
-In this process, this extra LLM rewrite each labeled or query text and generates a long list of propositions, introducing more efficiency overhead compared to other models that use the LLM solely for classification or keyword extraction, which typically generate only a few tokens.
-It is also worth noting that, although GORAG’s online indexing mechanism introduces a vast amount of knowledge into the weighted graph, we can still achieve comparable or better time efficiency with other baselines.
-4.3. Zero-shot Experiments
-To further study GORAG’s 0-shot ability, we compare GORAG’s performance with some widely applied open-source LLM models without RAG approaches.
-We conducted experiments on the WOS dataset by providing only the label names, without any labeled data to GORAG for creating the graph, the result is shown in Table 4.
-In the 0-shot setting, GORAG first generates label descriptions based on the label names, then extracts keywords from these descriptions, without using any information from labeled texts.
-Note that the Entailment model cannot be applied in this setting, as they require labeled data for training.
-For other RAG-based baselines, they are considered equivalent to their backbone LLMs in the 0-shot setting due to a lack of external information sources.
-As shown by the results, GORAG achieves comparable performance with other open-source LLM models in the first round and outperforms them in all subsequent rounds.
-This is because, when the number of target labels is small, the benefit of compressing the target label set is not as pronounced.
-However, as the number of labels increases, the importance of filtering the label set becomes more significant, allowing GORAG to consistently outperform other open-source LLM models.
-Table 4. Zero-shot experiment result on WOS dataset.
-Model Round
-R1 R2 R3 R4
-Qwen2-7B (Yang et al., 2024) 0.3322 0.1790 0.1451 0.1444
-Mistral0.3-7B (Jiang et al., 2023) 0.1436 0.0540 0.0336 0.0270
-LLaMA3-8B (Touvron et al., 2023) 0.3351 0.1614 0.1161 0.0930
-GORAG 0.3305 0.2567 0.2230 0.2102
-4.4. Generalization Evaluation
-To evaluate the generalization ability of GORAG, we conducted additional experiments on the 6 and 8 round split versions of the WOS dataset under 1-shot and 5-shot settings, as the WOS dataset contains more testing data and labels than Reuters dataset, and more complex texts than the IFS-Rel dataset.
-In Figure 5, the results for rounds 1 to 4 are obtained from the 4-round split version, rounds 5 and 6 are obtained from the 6-round split version, and rounds 7 and 8 are obtained from the 8-round split version of the WOS dataset.
-This setting can ensure a sufficient amount of testing data at each round, As illustrated in Figure 5, the classification accuracy of GraphRAG and NaiveRAG drops significantly from round 1 to 8, while GORAG maintains competitive classification accuracy as the number of rounds increases through 1 to 8, with both 1 and 5-shot setting, demonstrating the importance of a more precise retrieval.
-Table 5. Ablation experiments on WOS for GORAG’s variants.
-Dataset Model Round
-R1 R2 R3 R4
-WOS GORAG offline 0.3063 0.2302 0.2455 0.2156
-GORAG unit 0.4706 0.4394 0.4407 0.3899
-GORAG keyword 0.4746 0.4606 0.4455 0.4030
-GORAG 0.4862 0.4649 0.4814 0.4210
-Reuters GORAG offline 0.0000 0.1000 0.0917 0.1032
-GORAG unit 0.0000 0.0625 0.1667 0.0795
-GORAG keyword 0.0000 0.1000 0.1167 0.1438
-GORAG 0.0875 0.1750 0.1667 0.1667
+To evaluate the efficiency of GORAG, we compare the total time cost of GORAG and other baselines on the 4 round split version of WOS and CAD datasets with Qwen2.5-7B backbone, the results are shown in Figure 3(a) and 3(b).
+Although GORAG’s online indexing mechanism introduces a vast amount of knowledge into the weighted graph, we can still achieve comparable or better time efficiency with other baselines.
+We also list the token usage comparison between GORAG and other baselines in Table 4, and the graph’s memory usage and incremental trends in Table 3, where GORAG consumes fewer tokens than current RAG baselines, and only introduce less than 50KB memory cost compared with solely applying LLMs.
 4.3. Ablation Study
-To analyze GORAG’s superior performance, we conduct an extensive ablation study in the 1-shot setting, as RAG models show minimal performance changes across shot settings.
+To further analyze GORAG’s performance, we conduct an extensive ablation study under 1-shot setting with Qwen2.5-7B backbone, as RAG models show similar performance across shot settings.
 Specifically, we examine how different GORAG components impact classification performance using its variants.
@@
 GORAG keyword only uses the keyword extracted from the text to create LLM classification input, rather than the whole text.
-As shown in Table 5, firstly, GORAG offline achieves the worst accuracy, this is because it lacks a comprehensive retrieval source, making its generated candidate labels inaccurate, hence limits its performance.
+Table 4. Token cost on IFS-Rel dataset with 1-shot setting.
+Model AVG Token per Query AVG Cost per Query (USD)
+GPT-4o 634.6 \$0.005
+LongLLMLingua 2771.3 \$0.021
+NaiveRAG 4537.4 \$0.034
+LightRAG 7866.3 \$0.059
+GraphRAG 9296.3 \$0.070
+GORAG 557.5 \$0.004
+As shown in Table 5, firstly, GORAG offline achieves the worst accuracy, because it lacks a comprehensive retrieval source, making its generated candidate labels inaccurate, hence limits its performance.
 Secondly, GORAG unit outperforms GORAG offline, but still sub-optimal to GORAG, demonstrates the importance of modeling the varying importance between keywords and each label, which leads to accurate retrievals.
 Lastly, GORAG keywords demonstrates the important information for text classification are mostly text keywords, as it achieves comparable or better performance than other ablated baselines.
+Table 5. Ablation experiments on WOS for GORAG’s variants.
+Dataset Model Round
+R1 R2 R3 R4
+WOS GORAG offline&unit 0.2979 0.2287 0.2645 0.2249
+GORAG offline 0.3063 0.2302 0.2455 0.2156
+GORAG unit 0.4706 0.4394 0.4407 0.3899
+GORAG keyword 0.4746 0.4606 0.4455 0.4030
+GORAG 0.4862 0.4649 0.4814 0.4210
+4.4. More Generalization Evaluations
+4.4.1. Zero-shot Experiments
+To further study GORAG’s 0-shot ability, we compare GORAG’s performance with the original powerful LLM GPT-4o (Hurst et al., 2024) without RAG approaches, and some RAG baselines that are applicable to the 0-shot scenario in Table 6.
+Due to cost constraints, we conducted experiments on the second largest IFS-Rel dataset by providing only the label names, without any labeled data to GORAG for creating the graph.
+In the 0-shot setting, we input the labels and their LLM generated descriptions as system prompt for each model, and GORAG would extract keywords from label descriptions, without using any information from labeled texts.
+Note that the Entailment model cannot be applied in this setting, as they require labeled data for training.
+For other RAG-based baselines, they are considered equivalent to their base LLMs in the 0-shot setting due to a lack of external information.
+According to Table 6, GORAG can still achieve better performance than all tested baselines.
+4.4.2. Round Generalization Evaluation
+To evaluate GORAG’s ability to generalize to a larger number of dynamic rounds, we conducted additional experiments on the 6 and 8 round split versions of the WOS dataset under 1-shot and 5-shot settings, as the WOS dataset contains more testing data and labels than CAD dataset, and more lengthy, complex texts than the IFS-Rel dataset.
+In Figure 4, the results for rounds 1 to 4 are obtained from the 4-round split version, rounds 5 and 6 are obtained from the 6-round split version, and rounds 7 and 8 are obtained from the 8-round split version of the WOS dataset.
+This setting can ensure a sufficient amount of testing data at each round, As illustrated in Figure 4, the classification accuracy of GraphRAG and NaiveRAG drops significantly from round 1 to 8, while GORAG maintains competitive classification accuracy as the number of rounds increases through 1 to 8, with both 1 and 5-shot setting, demonstrating the importance of a more precise retrieval.
+Table 6. 0-shot experiments for GORAG and RAG baselines on IFS-Rel dataset under 0-shot setting.
+Model R1 R2 R3 R4
+GPT-4o (Hurst et al., 2024) 0.9219 0.5109 0.4063 0.5936
+LongLLMLingua (Jiang et al., 2024) 0.9281 0.5297 0.4234 0.5806
+LightRAG (Guo et al., 2025) 0.6718 0.3047 0.2135 0.0977
+GraphRAG (Edge et al., 2024) 0.5625 0.1875 0.1458 0.0938
+GORAG 0.9328 0.5406 0.4375 0.6063
+(a) 1-shot setting.
+(b) 5-shot setting.
+Figure 4. Experiment result of WOS dataset under 1-shot and 5-shot settings for at most 8 rounds.
 5. Conclusion
-In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Text Classification (DFSTC) task.
-Extensive experiments on three datasets with different characteristics demonstrate the effectiveness of GORAG in classifying texts with only a limited number or even no labeled data.
-Additionally, GORAG shows its effectiveness in adapting to the dynamic updates of target labels by adaptively retrieving candidate labels to filter the large target label set in each update round.
-With extensive of ablation studies, we confirm that GORAG’s edge weighting, adaptive retrieval, and online indexing mechanisms contribute to its effectiveness.
+In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Social media Text Classification (DFSTC) task.
+Extensive experiments on hate speech and misinformation detection benchmarks with different characteristics demonstrate the effectiveness of GORAG in classifying social media texts with only a limited number or even no labeled data.
+GORAG shows its effectiveness in adapting to the dynamic updates of target labels by adaptively retrieving candidate labels to filter the large target label set in each update round.
 For future work, we aim to further enhance GORAG’s performance and explore its application in more general scenarios.
+6. Acknowledgements
+Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No.
+2023YFF0725100, National Science Foundation of China (NSFC) under Grant No.
+U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No.
+2024A0505040012, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Key Areas Special Project of Guangdong Provincial Universities 2024ZDZX1006, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023A03J0012, HKUST(GZ) - CMCC(Guangzhou Branch) Metaverse Joint Innovation Lab under Grant No.
+P00659, Hong Kong ITC ITF grant PRP/004/22FX, Zhujiang scholar program 2021JC02X170, HKUST-Webank joint research lab.
 Appendix A Notations
-Due to the space limitation, we put the summary table on important notations here.
+Due to the space limitations, we put the summary table on important notations here in Table 7.
 Table 7. Summary on Important Notations.
@@
 $w$ The edge weight.
-$r$ The round of label updation.
+$r$ The round of label updating.
+$\tau$ The keyword extraction threshold.
 $y\in\mathcal{Y}^{r}$ The target label at round $r$.
@@
 $\mathcal{G}^{r}(\mathcal{V},\mathcal{E})$ weighted graph with node set $\mathcal{V}$ and edge set $\mathcal{E}.$
-$\mathcal{Y}^{r}$, $\mathcal{Y}^{r}_{new}$ All target label and new target labels at round $r$.
-$\mathcal{D}^{r}$, $\mathcal{D}^{r}_{new}$ The labeled training text for all seen labels at round $r$.
-$\mathcal{T}^{r}$, $\mathcal{T}_{new}^{r}$ All seen texts at round $r$.
+$\mathcal{Y}^{r}$ New target labels at round $r$.
+$\mathcal{D}^{r}$ The labeled training text for all new labels at round $r$.
+$\mathcal{T}^{r}$ Unlabeled texts at round $r$.
 $\mathcal{T}_{test}^{r}$ The query text for at round $r$.
-$\mathcal{K}_{new}^{r}$ The extracted keyword set for new labels of round $r$.
+$\mathcal{K}^{r}$ The extracted keyword set for new labels of round $r$.
 $\mathcal{V}_{new}^{r}$ The set of new graph nodes to be added in round $r$.
@@
 $P_{u,v}$ The shortest path between node $u$ and node $v$.
-$\mathcal{V}_{exist}$ The keywords that exist in graph.
-$\mathcal{V}_{notexist}$ The keywords do not exist in graph.
+$\mathcal{V}^{t}_{exist}$ The keywords extracted from $t$ and exist in graph.
+$\mathcal{V}^{t}_{notexist}$ The keywords extracted from $t$ and not exist in graph.
 $\mathcal{E}_{oi}^{t}$ Edges added based on text $t$.
-Appendix B Complexity Analysis
-In this section, we analysis the time and space complexity of GORAG’s graph construction, retrieval and prediction procedure.
-We denote the maximum number of terms of the input text as $m_{t}$, the number of unique terms of training corpus $\mathcal{D}_{new}^{r}$ as $u$, the LLM’s maximum input token length as $m_{i}$, LLM’s its maximum extraction, generation, and classification token length as $m_{e}$, $m_{g}$, and $m_{c}$, respectively.
-B.1. Graph Construction Complexity
-Firstly, we analysis the time and space complexity of GORAG’s graph construction.
-For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e}))$; If label names are available, the time complexity of generat label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}))$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$ times, and merging graph requires
-$O(|\mathcal{Y}^{r}_{new}||\mathcal{Y}^{r-1}|)$. Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
-$O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e})+|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}+|\mathcal{Y}^{r-1}|)+|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$.
-For the space complexity, the graph is stored with the weighted adjency matrix, hence needs $O(|\mathcal{E}^{r}|)$ space; Storing the training corpus at the $r$-th round would need $O(u|\mathcal{D}_{new}^{r}|)$ space; Storing the representive keywords $\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}$ would need $O(|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$ space.
-Hence the total space complexity of GORAG’s graph construction at the $r$-th round would be:
-$O(|\mathcal{E}^{r}|+u|\mathcal{D}_{new}^{r}|+|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$.
-B.2. Retrieval and Classification Complexity
-The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (Mehlhorn, 1988b); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)$; The time complexity of the final classification by LLM is $O(m_{i}+m_{c})$.
-Hence, the total time complexity of GORAG’s adaptive retrieval and classification can be denoted as
-$O(|\mathcal{T}_{test}^{r}|(|\mathcal{E}^{r}|+|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|+m_{i}+m_{c})).$
-For the space complexity, $O(|\mathcal{E}^{r}|)$ to store the graph, and $O(u|\mathcal{T}_{test}^{r}|)$ space is needed to store the testing corpus.
-Hence, the total space complexity of GORAG’s Retrieval and Classification procedure is
-$O(|\mathcal{E}^{r}|+u|\mathcal{T}_{test}^{r}|).$
+Appendix B Proof
+Here we prove our Adaptive Candidate Label Generation problem can be reduced from the NP-hard Steiner Tree problem (Su et al., 2020).
+Theorem B.1.
+The Adaptive Candidate Label Generation problem is NP-hard.
+Proof.
+To demonstrate that the Adaptive Candidate Label Generation problem is NP-hard, we provide a simple reduction of our problem from the Steiner Tree problem.
+Since the Steiner Tree problem is already proven to be NP-hard (Su et al., 2020), we show that there is a solution for the Steiner Tree problem if and only if there is a solution for our problem.
+Firstly, given a solution $S$ for our problem, we can construct a Steiner Tree by generating a minimum spanning tree for all nodes in $S$ on graph $\mathcal{G}^{r}$ then connecting all nodes from $\mathcal{V}_{exist}^{t}$ to their closest neighbor nodes in $S$.
+Secondly, any Steiner Tree $ST$ that contains any node $y\in\mathcal{V}^{r}_{t}$ is also a solution of our problem with $|\{y\in\mathcal{V}^{r}_{t}\}|$ nodes.
+Thus, we prove that the adaptive candidate type generation problem is NP-hard.
+∎
 Appendix C Pseudo Code of GORAG
 In this section, we provide the pseudo code of GORAG’s graph construction and adaptive candidate type generation algorithm in 1 and 2, respectively.
-Appendix D Hardness Proof of The Adaptive Candidate Type Generation problem
-Theorem D.1.
-The Adaptive Candidate Type Generation problem is NP-hard.
-Proof.
-To demonstrate that the Adaptive Candidate Type Generation problem is NP-hard, we provide a simple reduction of our problem from the Steiner Tree problem.
-Since the Steiner Tree problem is already proven to be NP-hard (Su et al., 2020), we show that there is a solution for the Steiner Tree problem if and only if there is a solution for our problem.
-Firstly, given a solution $S$ for our problem, we can construct a Steiner Tree by generating a minimum spanning tree for all nodes in $S$ on graph $\mathcal{G}^{r}$ then connecting all nodes from $\mathcal{V}_{exist}^{t}$ to their closest neighbor nodes in $S$.
-Secondly, any Steiner Tree $ST$ that contains any node $y\in\mathcal{V}^{r}_{t}$ is also a solution of our problem with |$|\{y\in\mathcal{V}^{r}_{t}\}|$ nodes.
-Thus, we prove that the adaptive candidate type generation problem is NP-hard.
-∎
+Appendix D Complexity Analysis
+Firstly, we analyze the time and space complexity of GORAG’s graph construction.
+For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}^{r}|)$; If label names are available, the time complexity of generating label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}|)$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}||\mathcal{D}^{r}|)$ times, and merging graph requires $O(|\mathcal{Y}^{r}||\mathcal{Y}^{r-1}|)$.
+Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
+$O(|\mathcal{Y}^{r}||\mathcal{Y}^{r-1}|+|\mathcal{V}_{new}^{r}||\mathcal{D}^{r}|)$.
+[…]
+[…]
+The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (MEHLHORN1988125); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}||\mathcal{T}_{test}^{r}|$; Hence, the total time complexity of GORAG’s adaptive retrieval can be denoted as
+$O(|\mathcal{T}_{test}^{r}|(|\mathcal{E}^{r}|+|\mathcal{V}_{notexist}^{r}||\mathcal{T}_{test}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)).$
+For the space complexity, $O(|\mathcal{E}^{r}|)$ to store the graph, and $O(|\mathcal{T}_{test}^{r}|)$ space is needed to store the testing corpus.
+Hence, the total space complexity of GORAG’s Retrieval and Classification procedure is
+$O(|\mathcal{E}^{r}|+|\mathcal{T}_{test}^{r}|).$
 Algorithm 1 Graph Construction Algorithm of GORAG at round $r$
-0:  $\mathcal{D}_{new}^{r}$: Training text set at the $r$-th round and the provided label $y_{i}\in\mathcal{Y}_{new}^{r}$ for each training text $t_{j}$; $p_{extraction},\ p_{gen}$: The extraction and description generation instruction prompt; $\mathcal{G}^{r-1}$: Weighted graph of the previous round $r-1$.
-0:  $\mathcal{G}^{r}$: Constructed weighted graph at the $r$-th round.
+[…]
+[…]
 1: Let $\mathcal{V}_{text}^{r}=\emptyset$, $\mathcal{V}_{label}^{r}=\emptyset$, $\mathcal{E}_{new}^{r}=\emptyset$
-2:  for each text $t\in\mathcal{T}_{new}^{r}$ do
-3:     $\mathcal{V}_{text}^{r}=\mathcal{V}_{text}^{r}\cup LLM(p_{extract},t)\cup y_{t}$
+2:for each text $t\in\mathcal{T}^{r}$ do
+3:  $\mathcal{V}_{text}^{r}=\mathcal{V}_{text}^{r}\cup KeywordExtract(t)\cup y_{t}$
 4:end for
 5:if label names are available then
-6:     for each label $y_{i}\in\mathcal{Y}_{new}^{r}$ do
-7:        Get label description $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r})$.
-8:        $\mathcal{V}_{label}^{r}=LLM(p_{extract},\mathcal{K}_{y_{i}})\cup\mathcal{V}_{label}^{r}$
-9:     end for
-10:  end if
-11:  Let $\mathcal{V}_{new}^{r}=\mathcal{V}_{text}^{r}\cup\mathcal{Y}^{r}_{new}\cup\mathcal{V}_{label}^{r}$ be the new node set at the $r$-th round.
-12:  for each node $v\in\mathcal{V}_{new}^{r}$ do
-13:     $\mathcal{E}_{new}^{r}=\mathcal{E}_{new}^{r}\cup e_{v,y_{i}}$
-14:     $\mathcal{W}_{new}^{r}=\mathcal{W}_{new}^{r}\cup w_{v,y_{i}}^{r}$
-15:  end for
-16:  Let $\mathcal{G}^{r}_{new}=(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r})$ be the newly constructed graph for at the $r$-th round.
-17:  Let $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1})$ be the final constructed graph at the $r$-th round.
-18:  return  $\mathcal{G}^{r}$
+6:  for each label $y_{i}\in\mathcal{Y}^{r}$ do
+7:   Get some label related keywords $\mathcal{V}_{label}^{r}=LLM(p_{gen},\mathcal{D}^{r})\cup\mathcal{V}_{label}^{r}$.
+8:  end for
+9:end if
+10: Let $\mathcal{V}_{new}^{r}=\mathcal{V}_{text}^{r}\cup\mathcal{Y}^{r}_{new}\cup\mathcal{V}_{label}^{r}$ be the new node set at the $r$-th round.
+11:for each node $v\in\mathcal{V}_{new}^{r}$ do
+12:  $\mathcal{E}_{new}^{r}=\mathcal{E}_{new}^{r}\cup e_{v,y_{i}}$
+13:  $\mathcal{W}_{new}^{r}=\mathcal{W}_{new}^{r}\cup w_{v,y_{i}}^{r}$
+14:end for
+15: Let $\mathcal{G}^{r}_{new}=(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r})$ be the newly constructed graph for at the $r$-th round.
+16: Let $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1})$ be the final constructed graph at the $r$-th round.
+17:return $\mathcal{G}^{r}$
 Appendix E Experiment Details
-E.1. Further introduction of chosen baselines
+E.1. Introduction of chosen baselines
+As shown in Table 2, we compare GORAG’s performance with 7 baselines from 3 technical categories for few-shot experiments, they are NN-based Entailment (Xia et al., 2021), Long Context-based NaiveRAG (Gao et al., 2023), LongLLMLingua (Jiang et al., 2024), and Graph-based GraphRAG (Edge et al., 2024), LightRAG (Guo et al., 2025), and HippoRAG2 (gutiérrez2025ragmemory).
 NN-based Models
   • Entailment (Xia et al., 2021): Entailment concatenates the text data with each of the label names to form multiple entailment pairs with one text sample, hence increasing the number of training data and enhance its finetuning of a RoBERTa PLM (Liu, 2019).
-Entailment then carry out classification based on these entailment pairs and convert the text classification task into a binary classification task, for detecting whether the entailment pair formed is correct.
-When all entailment pairs of a particular text are considered to be incorrect, Entailment will directly give up the classification of this text.
-In this paper, we consider this classification as a hallucination.
 Long Context RAG Models
-  • NaiveRAG (Gao et al., 2023): NaiveRAG acts as a foundational baseline of current RAG models.
-When indexing, it stores text segments of the labeled texts in a vector database using text embeddings.
-When querying, NaiveRAG generates query texts’ vectorized representations to retrieve side information based on the highest similarity in their embeddings.
-  • Propositionizer (Chen et al., 2023): Propositionizer applies a fine-tuned LLM to convert side information into atomic expressions, namely propositions, to facilitate more fine-grained information retrieval than NaiveRAG.
-Compression RAG Models
-  • LongLLMLingua (Jiang et al., 2024a): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
-  • GraphRAG (Edge et al., 2024): GraphRAG is a graph-Based RAG model that employs the LLM to extract texts’ entities and relations, which are then represented as nodes and edges in the weighted graph.
-GraphRAG then aggregates nodes into communities, and generates a comprehensive community report to encapsulate global information from texts.
-  • LightRAG (Guo et al., 2024): LightRAG skips the GraphRAG’s formulation of graph communities and directly retrieve nodes or pathes from created graphs.
-It calculates the embedding similarity of text extracted entities with graph nodes, to achieve a one-to-one mapping from keywords to graph nodes and retrieve all triples involved these nodes.
+  • NaiveRAG (Gao et al., 2023): It acts as a foundational baseline of current RAG models.
+When indexing, it stores text segments in a vector database using text embeddings.
+When querying, retrieves the side information based on their vector similarity to the query.
+  • LongLLMLingua (Jiang et al., 2024): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
+Graph-based RAG Models
+  • GraphRAG (Edge et al., 2024): It employs LLM for graph construction.
+When retrieval, it aggregates nodes into communities w.r.t.
+the query, and generates community reports to encapsulate global information from texts.
+  • LightRAG (Guo et al., 2025): It calculates the vector similarity between query extracted entities and graph nodes, achieving a one-to-one mapping from entities to graph nodes, then retrieve these nodes’ graph neighborhood.
+  • HippoRAG2 (gutiérrez2025ragmemory): It employs LLM for graph construction, and also add the text chunks as nodes in the graph.
+As a result, it directly retrieves Top-k text chunks based on their respective nodes’ PPR score.
 Algorithm 2 Adaptive Candidate Type Generation Algorithm
-0:  $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$: The constructed weighted graph; $\mathcal{V}_{exist}^{t}\subseteq\mathcal{V}^{r}$: A set of keyword nodes extracted from text $t$ and can be mapped to graph $\mathcal{G}$; $\mathcal{Y}^{r}$: The target label set at the $r$-th round.
-0:  $\hat{\mathcal{Y}}^{r}_{t}$: The candidate type retrieved for text $t$.
+[…]
+[…]
 1: Compute a minimum spanning tree $MST$ of the graph $\mathcal{G}$.
-2:  Let $\mathcal{V}_{term}=\mathcal{V}_{exist}^{t}\cap\mathcal{V}(MST)$ be the terminal nodes in $MST$
+2: Let $\mathcal{V}_{term}=\mathcal{V}_{exist}^{t}\cap\mathcal{V}(MST)$ be the terminal nodes in MSTMST
 3: Construct a weighted auxiliary graph $H$: $\mathcal{V}(H)=\mathcal{V}_{term}$
 4:for each pair of terminals $u,v\in\mathcal{V}_{term}$ do
-5:     Find the shortest path $P_{uv}$ in $MST$ from $u$ to $v$
+5:  Find the shortest path $P_{uv}$ in MSTMST from $u$ to $v$
 6:  Let $w^{H}_{u,v}=\sum_{e\in P_{uv}}w_{e}$
@@
 10:for each edge $e_{u,v}\in MST^{\prime}$ do
-11:     Add the shortest path $P_{uv}$ in $MST$ to $ST$
+11:  Add the shortest path $P_{uv}$ in $MST$ to STST
 12:end for
@@
 14:return $\hat{\mathcal{Y}}^{r}_{t}$
-E.2. Hyperparameter and Hardware Settings
-In each round of our experiments, we handle NN and RAG-based models differently for training and indexing.
-For NN-based models, we train them using all labeled data from the current round and all previous rounds.
-For RAG-based models, we index the labeled data of the current round to the information source from the previous round.
-The RAG-based model is initialized from scratch only in the first round.
-After training or indexing in each round, we test the models using the testing data from the current round and all previous rounds.
-For GORAG, which employs an online indexing mechanism, we first test its performance with online indexing on the current round’s testing data.
-We then test its performance on the testing data from all previous rounds without online indexing.
-This approach allows us to study the effect of online indexing on the performance of later round data when applied to previous round data.
-For the hyperparameter settings, we train the Entailment model with the RoBERTa-large PLM for 5 epochs with a batch size of 16, a learning rate of $1\times 10^{-6}$, and the Adam optimizer (Kingma, 2014), following the exact setting in Entailment’s original paper.
-For LongLLMLingua, we test the compression rate within ${0.75,0.8,0.85}$ and select 0.8, as it achieves the best overall classification accuracy.
-For GraphRAG and LightRAG, we use their local search mode, as it achieves the highest classification accuracy on both the WOS and Reuters datasets.
-For GORAG and all RAG-based baselines, we use LLaMA3 (Touvron et al., 2023) as the LLM backbone unless otherwise specified.
+E.2. Hardware Settings
 All experiments are conducted on an Intel(R) Xeon(R) Gold 5220R @ 2.20GHz CPU and a single NVIDIA A100-SXM4-40GB GPU.
-E.3. Further Analysis based on Precision and F1
-According to Table 8 and Table 2, although Entailment achieves the second best accuracy than GORAG on WOS dataset, it has a high hallucination rate.
-This is because the ground truth label of Entailment’s entailment pairs are greatly unbalanced, hence make the model easy to overfit, and tend to predict all entailment pairs of a particular text as incorrect under few-shot settings.
-(a) GORAG’s candidate selection algorithm returns candidate labels adaptively based on keywords extracted from the tesing texts.
-(b) GORAG’s candidate selection algorithm only returns the ground truth label of the query text.
-Figure 6. Two case of GORAG successfully generate candidate labels to help text classification on WOS dataset.
-E.4. GORAG with different backbone LLMs
-Table 7. Evaluation on different LLM backbones.
-Dataset Model Round
-R1 R2 R3 R4
-WOS GORAG LLaMA3 0.4862 0.4649 0.4814 0.4210
-GORAG Qwen2.5 0.5101 0.4839 0.4823 0.4235
-In Table 7, we present the results of GORAG with different backbone LLMs, we test GORAG with more advanced Qwen2.5-7B (Qwen Team, 2024), the result demonstates the GORAG’s performance can be further improved with more advanced LLM backbones.
-E.5. Case Study
-In this section, to illustrate the strength of GORAG’s adaptive retrieval and candidate label generation procedure, we dig into two testing cases select from WOS dataset where GORAG’s adaptive retrieval helps to reduce the LLM input length and benefit text classification.
-As shown in Figure 6(a), for the case whose ground truth label is Algorithm Design, the candidate label retrieved by GORAG are Computer programming and Algorithm design.
-As a result, GORAG’s adaptive retrieval algorithm successfully filter the target label set to only contains two candidate labels, and successfully cut down the target label number from at most over 100 to only 2 candidate labels that it consided as the possible labels for the text, significantlly reduce the LLM’s input length and mitigate the lost-in-the-middle issue.
-On the other hand, for the case in Figure 6(b) whose ground truth label is Electrical Circuits, since the ground truth label name Electrical Circuits already exists in the text and being extracted as keyword, GORAG’s adaptive retrieval algorithm is almost certain to classify this text into class Electrical Circuits, hence GORAG’s adaptive retrieval and candidate type generation algorithm would only select the only candidate label retrieved from the constructed graph and input to the LLM.
-Table 8. The macro recall score and hallucination rate (Hall. Rate) on WOS dataset.
-Metrics Category Model Round
-R1 R2 R3 R4
-1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
-Recall NN-based Entailment 0.3502 0.3475 0.3908 0.3021 0.3415 0.3442 0.2327 0.2406 0.2432 0.1789 0.1792 0.1837
-Long Context RAG NaiveRAG 0.3324 0.3417 0.3398 0.1615 0.1652 0.1650 0.0943 0.1013 0.0872 0.0568 0.0805 0.0628
-Propositionizer 0.1032 0.1131 0.1114 0.0947 0.1210 0.1187 0.1352 0.1298 0.1254 0.1373 0.1345 0.1402
-Compression RAG LongLLMLingua 0.3139 0.3105 0.3258 0.1349 0.1388 0.1350 0.0862 0.0837 0.0872 0.0616 0.0610 0.0628
-GraphRAG 0.3791 0.3646 0.3732 0.1535 0.1519 0.1557 0.1009 0.1014 0.1015 0.0902 0.0940 0.0912
-LightRAG 0.3462 0.3441 0.3519 0.1544 0.1507 0.1466 0.1241 0.1257 0.1226 0.1129 0.1180 0.1101
-GORAG 0.4387 0.4659 0.4874 0.3074 0.3472 0.3483 0.2371 0.2513 0.2445 0.1960 0.1970 0.2013
-Hal. Rate NN-based Entailment 54.53% 58.96% 51.63% 44.76% 36.62% 37.76% 32.57% 28.14% 27.35% 29.68% 26.86% 24.13%
-Long Context RAG NaiveRAG 3.00% 2.32% 2.63% 6.45% 6.16% 6.28% 6.93% 7.46% 7.52% 9.96% 9.50% 9.83%
-Propositionizer 6.85% 5.88% 6.97% 7.40% 6.52% 6.09% 5.97% 6.44% 6.30% 5.78% 6.11% 5.52%
-Compression-RAG LongLLMLingua 6.50% 6.29% 5.63% 16.45% 16.05% 16.28% 19.63% 20.47% 20.52% 21.96% 21.35% 21.83%
-GraphRAG 2.40% 2.23% 2.07% 6.03% 5.97% 6.18% 7.20% 7.58% 7.03% 6.90% 6.39% 6.49%
-LightRAG 1.70% 1.24% 1.37% 2.59% 2.78% 3.21% 4.66% 4.55% 4.58% 5.89% 5.80% 5.73%
-GORAG 0.66% 1.16% 0.74% 0.30% 0.68% 0.65% 0.41% 0.57% 0.68% 0.40% 0.43% 0.64%
-Table 9. The macro recall score and hallucination rate (Hall. Rate) on IFS-Rel dataset.
-Metrics Category Model Round
-R1 R2 R3 R4
-1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
-Recall NN-based Entailment 0.3191 0.6015 0.5515 0.2172 0.3043 0.3330 0.1225 0.1802 0.2081 0.1093 0.1267 0.1490
-Long Context RAG NaiveRAG 0.4039 0.4194 0.4292 0.2320 0.2009 0.2115 0.1147 0.1163 0.1097 0.0871 0.0895 0.0912
-Propositionizer 0.1053 0.1146 0.1104 0.1124 0.1201 0.1197 0.0824 0.1029 0.0902 0.0744 0.0717 0.0799
-Compression RAG LongLLMLingua 0.4426 0.4625 0.4044 0.1879 0.1559 0.2158 0.0947 0.0906 0.1113 0.0698 0.0822 0.0696
-GraphRAG 0.4264 0.4139 0.4208 0.2116 0.2008 0.2047 0.1213 0.1179 0.1185 0.0906 0.0857 0.0883
-LightRAG 0.4221 0.3985 0.4382 0.2344 0.2150 0.2429 0.1270 0.1313 0.1276 0.0902 0.0926 0.0918
-GORAG 0.4588 0.4681 0.5000 0.2348 0.2175 0.2405 0.1279 0.1295 0.1290 0.0950 0.0886 0.0919
-Hal. Rate NN-based Entailment 8.10% 15.63% 4.69% 26.41% 19.14% 6.25% 26.56% 17.86% 6.67% 6.48% 5.90% 3.28%
-Long Context RAG NaiveRAG 1.56% 4.25% 1.78% 3.05% 2.97% 3.13% 3.02% 3.70% 3.54% 4.02% 4.18% 3.63%
-Propositionizer 7.82% 7.34% 11.25% 5.78% 6.40% 6.72% 4.84% 6.09% 7.03% 8.13% 8.59% 6.72%
-Compression-RAG LongLLMLingua 3.59% 2.97% 6.56% 2.12% 6.23% 5.47% 6.36% 6.00% 3.50% 3.29% 5.65% 8.88%
-GraphRAG 4.25% 1.56% 1.25% 2.66% 3.98% 2.89% 3.39% 3.75% 3.80% 3.91% 3.24% 3.09%
-LightRAG 3.28% 2.66% 2.03% 2.73% 2.97% 2.27% 1.98% 1.56% 1.67% 2.46% 2.23% 2.30%
-GORAG 0.31% 0.63% 0.16% 1.78% 2.81% 2.58% 1.30% 0.78% 1.09% 1.48% 1.02% 1.25%
+E.3. Case Study
+In this section, we dig into a test case from the CAD social media abuse detection dataset to demonstrate the ability of GORAG to provide graph-grounded clues, to explain its classification result.
+As shown in Figure 5, the keyword entity feminists and sexual predators are extracted from the input unlabeled text and mapped to the graph.
+GORAG then generates a Minimum Cost Steiner Tree through these mapped entities, which also spans through the candidate label Insult.
+The structure and involved nodes of this generated Steiner Tree (edges in red) serve as an explicit clue, helping the LLM to classify the text correctly as Insulting Content.
+Figure 5. A case from the CAD dataset, where the generated Steiner Tree serves as an explicit clue, helping the LLM to classify the text correctly.